### v2

# Three Fundamental Questions in Modern Infinite-Domain Constraint Satisfaction

Abstract.
The Feder-Vardi dichotomy conjecture for Constraint Satisfaction Problems (CSPs) with finite templates, confirmed independently by Bulatov and Zhuk, has an extension to certain well-behaved infinite templates due to Bodirsky and Pinsker which remains wide open. We provide answers to three fundamental questions on the scope of the Bodirsky-Pinsker conjecture.
Our first two main results provide two simplifications of this scope, one of structural, and the other one of algebraic nature. The former simplification implies that the conjecture is equivalent to its restriction to templates without algebraicity, a crucial assumption in the most powerful classification methods. The latter yields that the higher-arity invariants of any template within its scope can be assumed to be essentially injective, and any algebraic condition characterizing any complexity class within the conjecture closed under Datalog reductions must be satisfiable by injections, thus lifting the mystery of the better applicability of certain conditions over others.
Our third main result uses the first one to show that any non-trivially tractable template within the scope serves, up to a Datalog-computable modification of it, as the witness of the tractability of a non-finitely tractable finite-domain Promise Constraint Satisfaction Problem (PCSP) by the so-called sandwich method. This generalizes a recent result of Mottet and provides a strong hitherto unknown connection between the Bodirsky-Pinsker conjecture and finite-domain PCSPs.
1. Introduction
1.1. The finite-domain CSP dichotomy theorem
Fixed-template Constraint Satisfaction Problems are computational problems parametrized by structures $\mathfrak{A}$ with a finite relational signature, called templates; they are denoted by $\mathrm{CSP}(\mathfrak{A})$ and ask whether a given structure $\mathfrak{J}$ with the same signature as $\mathfrak{A}$ admits a homomorphism to $\mathfrak{A}$. The general CSP framework is incredibly rich, in fact, it contains all decision problems up to polynomial-time equivalence [21]. For this reason, CSPs are typically studied under additional structural restrictions on the template $\mathfrak{A}$. The restriction that has received most attention is requiring the domain of $\mathfrak{A}$ to be finite, and the problems arising in this way are called finite-domain CSPs. Already in this setting, one obtains many well-known problems such as HORN-SAT, 2-SAT, 3-SAT, or 3-COLORING.
Consider the concrete example of 3-COLORING. This prominent NP-complete problem can be modeled as $\mathrm{CSP}(\mathfrak{K}_{3})$, where $\mathfrak{K}_{3}$ stands for the complete graph on three vertices: indeed, a homomorphism to $\mathfrak{K}_{3}$ from a given structure $\mathfrak{J}$ in the same signature (i.e. from a given graph $\mathfrak{J}$) is simply a mapping that does not map any two adjacent vertices of $\mathfrak{J}$ to the same vertex of $\mathfrak{K}_{3}$. A key observation in the theory of CSPs is that the NP-completeness of $\mathrm{CSP}(\mathfrak{K}_{3})$ can be extended to CSPs of other structures which can “simulate” $\mathfrak{K}_{3}$ over their domain or over a finite power thereof using primitive positive (pp) formulas. Enhancing this by the additional observation that homomorphically equivalent templates have equal CSPs, one arrives at the notion of pp-constructibility [9]. If $\mathfrak{A^{\prime}}$ is pp-constructible from $\mathfrak{A}$, then $\mathrm{CSP}(\mathfrak{A^{\prime}})$ is reducible to $\mathrm{CSP}(\mathfrak{A})$ in logarithmic space; moreover, this reduction can be formulated in the logic programming language Datalog. This uniform reduction between CSP templates is sufficiently powerful, as we will see below, to describe all NP-hardness amongst finite-domain CSPs.
In the early 2000s, the field of finite-domain constraint satisfaction quickly rose in fame, in particular due to the discovery of tight connections with universal algebra. The most basic of these connections is that two structures with identical domains have the same sets of pp-definable relations if and only if they have the same sets of polymorphisms [51, 50]. Here, a polymorphism of a relational structure $\mathfrak{A}$ is simply a homomorphism from a finite power of $\mathfrak{A}$ into $\mathfrak{A}$ itself; we denote by $\mathrm{Pol}(\mathfrak{A})$ the set of all polymorphisms of $\mathfrak{A}$, the polymorphism clone of $\mathfrak{A}$. Since taking expansions by primitive positive relations does not lead to an increase in complexity, the study of finite-domain CSPs is subsumed by the study of finite algebras. Over the past two decades, the link between universal algebra and CSPs was gradually refined. After an intermediate stop at pp-interpretations [36], today we know that also pp-constructibility between finite relational structures is fully encoded in their polymorphism clones [9, Theorem 1.3]: $\mathfrak{A^{\prime}}$ is pp-constructible from $\mathfrak{A}$ if and only if $\mathrm{Pol}(\mathfrak{A^{\prime}})$ satisfies all height-1 identities of $\mathrm{Pol}(\mathfrak{A})$.
[…]
Theorem 1.1 (Bulatov and Zhuk [35, 66]).
Let $\mathfrak{A}$ be a finite relational structure. Then either $\mathfrak{A}$ pp-constructs $\mathfrak{K}_{3}$ and $\mathrm{CSP}(\mathfrak{A})$ is NP-complete, or $\mathrm{Pol}(\mathfrak{A})$ satisfies the Siggers identity
$s(x,y,z,x,y,z)\approx s(y,z,x,z,x,y)$
and $\mathrm{CSP}(\mathfrak{A})$ is solvable in polynomial time.
1.2. The infinite-domain CSP tractability conjecture
Arguably, the two main reasons for the popularity of finite-domain CSPs over CSPs which require an infinite template are immediate containment of such problems in the class NP (a homomorphism can be guessed and verified in polynomial time) as well as the above-mentioned applicability of algebraic methods which had been developed independently of CSPs for decades. Yet, even within the class NP, finite-domain CSPs are of an extremely restricted kind: already simple problems such as ACYCLICITY of directed graphs (captured by the template $(\mathbb{Q};<)$), various natural coloring problems for graphs such as NO-MONOCHROMATIC-TRIANGLE, and more generally the model-checking problem for natural restrictions of existential second-order logic such as the logic MMSNP of Feder and Vardi [43], are beyond its primitive scope despite their containment in NP.
The quest for a CSP framework that includes such problems whilst staying within NP and allowing for an algebraic approach akin to the one for finite templates started with the popularization of $\omega$-categoricity by Bodirsky and Nešetřil [28] as a sufficient structural restriction ensuring the latter: for countable $\omega$-categorical structures, pp-definability is determined by their polymorphism clone and, as was subsequently shown, so are the general reduction of pp-interpretability [30] as well as the pp-constructibility of $\mathfrak{K}_{3}$ [9]. The $\omega$-categoricity of $\mathfrak{A}$ can be interpreted as $\mathfrak{A}$ being “finite modulo automorphisms”: more precisely, the action of its automorphism group $\mathrm{Aut}(\mathfrak{A})$ on $d$-tuples has only finitely many orbits for all finite $d\geq 1$. It is, however, a mathematical property with little computational bearing: the CSPs of $\omega$-categorical structures can be monstrous from this perspective [48, 47]. For this reason, and based on strong empirical evidence, Bodirsky and Pinsker [32] identified a proper subclass of countable $\omega$-categorical structures as a candidate for an algebraic complexity dichotomy extending the theorem of Bulatov and Zhuk which does not seem to suffer from this deficiency.
Their first requirement on $\mathfrak{A}$ is that the orbit under $\mathrm{Aut}(\mathfrak{A})$ of any $d$-tuple be determined by the relations that hold on it: we call a relational structure $\mathfrak{A}$ homogeneous if every isomorphism between its finite substructures extends to an automorphism of $\mathfrak{A}$. This gives an effective way of representing orbits, which are central to $\omega$-categoricity. They then additionally impose an effective way of determining whether a given finite substructure is contained in $\mathfrak{A}$, which places the CSP in NP: a relational structure $\mathfrak{A}$ over a finite signature $\tau$ is finitely bounded [54] if there exists a finite set $\mathcal{N}$ of finite $\tau$-structures (bounds) such that a finite $\tau$-structure embeds into $\mathfrak{A}$ if and only if it does not embed any member of $\mathcal{N}$; equivalently, the finite substructures are up to isomorphism precisely the finite models of some universal first-order sentence [15, Lemma 2.3.14] (see also [26, 63] for more details).
Example 1.2.
A standard example of a countable finitely bounded homogeneous structure is $(\mathbb{Q};<)$. Its finite substructures are all finite strict linear orders, which can be axiomatized by irreflexivity, totality, and transitivity:
$\forall x,y,z\big{(}\neg(x<x)\wedge(x<y\vee x=y\vee y<x)\wedge(x<y\wedge y<z\Rightarrow x<z)\big{)}.$
Morevoer, every isomorphism between two such finite substructures can be extended to an automorphism of $(\mathbb{Q};<)$, e.g. by a piecewise affine transformation.
Finite boundedness and homogeneity taken together are very strong an assumption. Bodirsky and Pinsker observed that for all practical purposes, in particular the applicability of polymorphisms to complexity as well as containment in NP, it is sufficient that the template $\mathfrak{A}$ be a first-order reduct of a structure $\mathfrak{B}$ enjoying these, i.e. first-order definable therein. This yields sufficient flexibility to model a huge class of computational problems which includes in particular the ones mentioned above. In fact, requiring $\mathfrak{A}$ to be a reduct of a finitely bounded homogeneous structure $\mathfrak{B}$, i.e. obtained from $\mathfrak{B}$ by forgetting relations, turns out to be equivalent [3, Proposition 7], and we shall find this approach convenient.
Example 1.3.
The CSP of the template $(\mathbb{Q};\operatorname{Betw})$, first-order defined by
$\operatorname{Betw}(x,y,z):\Leftrightarrow(x<y<z)\vee(z<y<x)$
in the structure $(\mathbb{Q};<)$, is the classical NP-complete betweenness problem from the 1970s [60].
The following formulation of the conjecture of Bodirsky and Pinsker takes into account later progress [9, 10, 11, 7].
Conjecture 1.4 (Bodirsky and Pinsker 2012 [32]).
Let $\mathfrak{A}$ be a reduct of a countable finitely bounded homogeneous structure $\mathfrak{B}$. Then exactly one of the following holds.
  • $\mathfrak{A}$ pp-constructs $\mathfrak{K}_{3}$ (and consequently, $\mathrm{CSP}(\mathfrak{A})$ is NP-complete);
  • $\mathfrak{A}$ does not pp-construct $\mathfrak{K}_{3}$; in this case $\mathrm{Pol}(\mathfrak{A})$ satisfies the pseudo-Siggers identity
    $\alpha\circ s(x,y,z,x,y,z)\approx\beta\circ s(y,z,x,z,x,y)$
    and $\mathrm{CSP}(\mathfrak{A})$ is polynomial-time solvable.
It is important to note that the open part of Conjecture 1.4 is only the consequence of polynomial-time tractability: the negation of the first item does yield the satisfaction of the pseudo-Siggers identity by a theorem due to Barto and Pinsker [11]. It is even equivalent to it for model-complete cores [7, Theorem 1.3]: an $\omega$-categorical model-complete core is a template $\mathfrak{A}$ whose endomorphisms preserve all orbits of $\mathrm{Aut}(\mathfrak{A})$; it exists for all CSPs with an $\omega$-categorical template and is unique up to isomorphism [13].
2. Three fundamental questions: our contributions
The present paper aims to answer three fundamental questions around Conjecture 1.4, of which the first concerns its scope; the second the algebraic invariants of templates therein; and the third its connection with the rapidly evolving field of (finite-domain) Promise Constraint Satisfaction Problems. More precisely, we consider the following questions:
  1. (1)
    Are there significant additional structural assumptions, perhaps of model-theoretic nature, that can be imposed onto the structures of Conjecture 1.4 without loss of generality, i.e. without affecting the truth of the conjecture?
  2. (2)
    Are there significant algebraic assumptions that can be imposed onto the polymorphisms of the structures of Conjecture 1.4 without loss of generality?
  3. (3)
    Are there algorithmic connections between CSPs from Conjecture 1.4 and finite-domain Promise Constraint Satisfaction Problems?
We provide affirmative answers to all three questions, as we shall describe in the following.
2.1. Algebraicity is irrelevant
Conjecture 1.4 has been confirmed for many subclasses. Mottet and Pinsker speak of first- and second-generation classifications [59]: those of the former kind can roughly be described as exhaustive case-by-case analyzes, using Ramsey theory, of the available polymorphisms for structures first-order definable in a fixed finitely bounded homogeneous structure $\mathfrak{B}$ – extensive complexity classifications such as those for temporal CSPs [22], Graph-SAT problems [29], Poset-SAT problems [53], and CSPs with templates first-order definable in arbitrary homogeneous graphs [25] were obtained this way. Second-generation classifications, on the other hand, take a more structured approach mimicking advanced algebraic methods for finite domains; they were employed, for example, to achieve classifications for the logic MMSNP [24], Hypergraph-SAT problems [58], and certain graph orientation problems [44, 12]. However, even the most advanced methods as described in [59] require additional abstract structural assumptions on the template.
One such assumption is that the template $\mathfrak{A}$, or even the structure $\mathfrak{B}$ in which it is first-order defined, has no algebraicity. It is present virtually everywhere in the infinite-domain CSP literature [24]; and in particular an important prerequisite in most of the general results of the theory of smooth approximations developed by Mottet and Pinsker [59], as well as in all results of Bodirsky and Greiner [19, 20] about CSPs of generic superpositions of theories. We say that a structure $\mathfrak{A}$ has no algebraicity (in the group-theoretic sense) if, for every $k\geq 1$ and every tuple $\bar{a}\in A^{k}$, the automorphisms of $\mathfrak{A}$ which stabilize $\bar{a}$ do not stabilize any $a^{\prime}\notin\bar{a}$. In $\omega$-categorical structures this is the case if and only if, for every $k\geq 1$, every tuple $\bar{a}\in A^{k}$, and every $a^{\prime}\notin\bar{a}$, it is not possible to first-order define the unary relation $\{a^{\prime}\}$ using $\bar{a}$ as parameters; in other words, any non-trivial first-order property relative to $\bar{a}$ is always satisfied by infinitely many elements, which implies that some arguments become easier compared to finite structures (whereas naturally, many other arguments become harder). We will show that the following strengthening of this property (for model-complete cores, see Proposition 4.1 in Section 4.1) can be assumed when resolving Conjecture 1.4: a structure $\mathfrak{A}$ is CSP-injective if every finite structure that homomorphically maps to $\mathfrak{A}$ also does so injectively; in other words, if there is a solution to an instance of $\mathrm{CSP}(\mathfrak{A})$, then there is also an injective one. CSP-injectivity played an essential role in the universal-algebraic proof of the complexity dichotomy for Monotone Monadic SNP [24] and, more recently, certain finitely bounded homogeneous uniform hypergraphs [58].
Example 2.1.
The template $(\mathbb{Q};<)$ from Example 1.2 is CSP-injective because a digraph maps homomorphically to $(\mathbb{Q};<)$ if and only if it is acyclic; in this case there is always an injective homomorphism. It is a model-complete core, and hence has no algebraicity by Proposition 4.1 in Section 4.1. Its expansion $(\mathbb{Q};<)_{I_{4}}$ by the relation
$I_{4}\coloneqq\{(x,y,u,v)\mid x=y\Rightarrow u=v\}$
has the same automorphisms, and hence no algebraicity; it is no more CSP-injective by Propositions 4.2 and 4.3. The equivalence relation with countably many classes of size 2 has algebraicity: any automorphism stabilizing some element $a$ must also stabilize the other element of its class.
One of the properties preserved by our construction enforcing CSP-injectivity in Theorem 2.2 below is that of the structure $\mathfrak{B}$ being Ramsey. Although of central importance in classification methods [31, 61], its precise definition is not essential to the present paper and therefore omitted; we refer the reader to [49] for details about structural Ramsey theory.
Throughout the present paper, the notation $\mathfrak{A}\wr\mathfrak{B}$ stands for the relational wreath product of two structures $\mathfrak{A}$ and $\mathfrak{B}$, introduced formally in Section 4.2, and $\mathfrak{A}_{R}$ stands for the expansion of a structure $\mathfrak{A}$ by a relation $R$ defined over its domain.
Theorem 2.2 (No Algebraicity).
Let $\mathfrak{A}$ be a non-trivial reduct of a countable structure $\mathfrak{B}$ over a finite relational signature. Then $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ is a non-trivial CSP-injective reduct of $\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq}$, which is a countably infinite structure without algebraicity over a finite relational signature, and $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$ and $\mathrm{CSP}(\mathfrak{A})$ are Datalog-interreducible. Moreover:
  1. (1)
    If $\mathfrak{B}$ is $\omega$-categorical, then $\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq}$ is $\omega$-categorical as well. In this case:
    • $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ is a model-complete core if and only if $\mathfrak{A}$ is a model-complete core.
  2. (2)
    If $\mathfrak{B}$ is homogeneous, then $\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq}$ is homogeneous as well. In this case:
    • $\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq}$ is Ramsey if and only if $\mathfrak{B}$ is Ramsey.
  3. (3)
    If $\mathfrak{B}$ is finitely bounded, then $\smash{\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq}}$ is finitely bounded as well.
  4. (4)
    If the number of orbits of $d$-tuples of $\mathfrak{A}$ has growth slower than $2^{2^{d}}$, then:
    • $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ pp-constructs $\mathfrak{K}_{3}$ if and only if $\mathfrak{A}$ pp-constructs $\mathfrak{K}_{3}$.
It is not hard to see that item 4 of Theorem 2.2 applies to reducts of homogeneous structures in a finite relational signature [54], in particular to all structures in the scope of Conjecture 1.4.
2.2. Polymorphisms are injective
Comparing Theorem 1.1 with Conjecture 1.4, one notices the replacement of the Siggers identity by its (weaker) pseudo-variant. This is not just an inconvenience arising from the proof of Barto and Pinsker in [11] (or the more recent proof in [5]), but there are examples showing the necessity to weaken the condition.
Example 2.3.
The CSP of $(\mathbb{Q};<)_{I_{4}}$ from Example 2.1 is solvable in Datalog (Section 4.1). Since $\mathrm{CSP}(\mathfrak{K}_{3})$ is not, and pp-constructions provide Datalog-reductions [1], it follows that $(\mathbb{Q};<)_{I_{4}}$ does not pp-construct $\mathfrak{K}_{3}$. Hence its polymorphisms satisfy the pseudo-Siggers identity by the above-mentioned theorem of Barto and Pinsker. However, any polymorphism of $I_{4}$ is injective up to dummy variables [16], and hence cannot satisfy the Siggers identity.
We say that a template is Pol-injective if all of its polymorphisms are essentially injective, i.e. injective up to dummy variables – this is the case if and only if the relation $I_{4}$ is invariant under its polymorphisms. As it turns out, the relation $I_{4}$ can be added to the templates constructed in Theorem 2.2 at no computational cost (up to Datalog reductions). It thus follows that any algebraic condition on polymorphisms capturing a complexity class closed under Datalog-reductions for CSPs of structures within the scope of Conjecture 1.4 must necessarily be satisfiable by essentially injective operations. Until now, this fact was understood by experts on an intuitive level, but our result provides a rigorous confirmation.
Theorem 2.4.
The structures $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ and $\smash{\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq,I_{4}}}$ enjoy all properties of $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ and $\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq}$ from Theorem 2.2 except that CSP-injectivity is replaced by Pol-injectivity.
For finite structures the only essentially injective polymorphisms are at most unary up to dummy variables. It is therefore of no surprise that so few universal-algebraic conditions for CSPs from the finite have been successfully lifted to the infinite. One prominent example which can actually be lifted is the satisfiability of quasi near-unanimity identities, which captures bounded strict width both in the finite and in the infinite [17]; bounded strict width corresponds to solvability in a particular proper fragment of Datalog that is not closed under the Datalog-reductions in Theorem 2.2. Examples of conditions for polymorphism clones that can be satisfied by essentially injective operations are any pseudo height-1 identities (e.g., the pseudo-Siggers identity), but also the dissected near-unanimity identities [47] which were proposed in [33] as a possible candidate for solvability of CSPs in fixed-point logic.
Marimon and Pinsker have recently provided a negative answer, within the scope of Conjecture 1.4, to a question of Bodirsky [15, Question 14.2.6(27)] asking whether every $\omega$-categorical CSP template without algebraicity which is solvable in Datalog has a binary injective polymorphism [55, Corollary 8.10]. We contrast this negative result with a corollary to Theorem 2.4 that implies that up to Datalog reductions the answer is positive.
Corollary 2.5.
Let $\mathfrak{A}$ be a non-trivial reduct of a finitely bounded homogeneous structure $\mathfrak{B}$. If $\mathfrak{A}$ does not pp-construct $\mathfrak{K}_{3}$, then $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ has a binary injective polymorphism.
2.3. Everything is a cheese
One of the currently most vibrant branches of research in constraint satisfaction are Promise Constraint Satisfaction Problems (PCSPs). For two relational structures $\mathfrak{S}_{1}$ and $\mathfrak{S}_{2}$ such that $\mathfrak{S}_{1}$ maps homomorphically to $\mathfrak{S}_{2}$, the problem $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ asks to decide whether a given finite structure $\mathfrak{J}$ in the same signature as $\mathfrak{S}_{1}$ and $\mathfrak{S}_{2}$ maps homomorphically to $\mathfrak{S}_{1}$ or does not even map homomorphically to $\mathfrak{S}_{2}$; the instance $\mathfrak{J}$ is promised to satisfy one of those two cases. Thus, the problem $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ asks to compare $\mathrm{CSP}(\mathfrak{S}_{1})$ to a relaxation $\mathrm{CSP}(\mathfrak{S}_{2})$ of it, and since $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{1})=\mathrm{CSP}(\mathfrak{S}_{1})$, promise constraint satisfaction problems generalize the CSP framework. This generalization is vast, and contains many well-known computational problems such as APPROXIMATE GRAPH COLORING [45], $(2+\varepsilon)$-SAT [2], and HYPERGRAPH COLORING [40]. Although formally not necessary, research has focused on finite templates $(\mathfrak{S}_{1},\mathfrak{S}_{2})$: this is justified for example by the fact that even for structures over a Boolean domain, there is no complete complexity classification yet. Hence, the two extensions considered here of finite-domain CSPs to $\omega$-categorical templates, or alternatively to PCSPs, are somewhat orthogonal, and it is natural to wonder whether there exist any connections, in particular of algorithmic nature.
One potential link is the use of $\omega$-categorical CSP templates as cheeses in the sandwich method to prove tractability of PCSPs, as follows. Note that if $\mathfrak{A}$ is a structure such that $\mathfrak{S}_{1}$ maps homomorphically to $\mathfrak{A}$ and $\mathfrak{A}$ maps homomorphically to $\mathfrak{S}_{2}$, then $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ trivially reduces to $\mathrm{CSP}(\mathfrak{A})$; we call $\mathfrak{A}$ a cheese sandwiched by the template $(\mathfrak{S}_{1},\mathfrak{S}_{2})$. This simple observation is more powerful than it might first appear: in fact, so far, every known tractable (finite-domain) PCSP can be reduced to the tractable CSP of some cheese structure via the sandwich method [56, 39]. On the other hand, Barto [4] provided an example of a PCSP template sandwiching an infinite polynomial-time tractable cheese while not being finitely tractable, i.e. such that every finite-domain cheese has an NP-complete CSP [4]. Since Barto’s cheese is not $\omega$-categorical, he raised the question whether such an example could also be constructed with an $\omega$-categorical cheese ([4, Question IV.1]). Subsequently the question arose, and was promoted in particular by Zhuk at the CSP World Congress 2023, whether structures within the scope of Conjecture 1.4 could ever serve as tractable cheeses in the absence of a finite one.
The first two examples of tractable $\omega$-categorical cheeses for non-finitely tractable finite-domain PCSPs were recently obtained by Mottet [57]; his structures fall within the scope of Conjecture 1.4. In the present paper, we strengthen this result by creating non-finitely tractable finite-domain PCSPs from virtually every structure $\mathfrak{A}$ within the scope of Conjecture 1.4. Our construction consists of two steps. First, we show that such a construction is possible in general under the additional assumption that $\mathfrak{A}$ is equipped with an inequality predicate and that it is a reduct of a structure $\mathfrak{B}$ which is linearly ordered. We then combine this construction with Theorem 2.2. Modulo an extra step consisting of taking generic superpositions (Proposition 5.3) of structures with $(\mathbb{Q};<)$ (for which we need the property of no algebraicity), this finishes our proof.
Theorem 2.6 (Datalog-Approximability of Infinite-Domain CSPs).
For every non-trivial $\tau$-reduct $\mathfrak{A}$ of a countable finitely bounded homogeneous structure $\mathfrak{B}$ there exists a $\tau^{\prime}$-reduct $\mathfrak{A}^{\prime}$ of a countable finitely bounded homogeneous structure $\mathfrak{B}^{\prime}$ and Datalog-interpretations $\mathcal{J},\mathcal{J}^{\prime}$ from $\tau$-structures to $\tau^{\prime}$-structures and vice versa such that, for every $n\in\mathbb{N}$, there exists a finite PCSP template $(\mathfrak{S}_{1},\mathfrak{S}_{2})$ with the following properties:
  1. (1)
    $\mathfrak{A}$ pp-constructs $\mathfrak{K}_{3}$ if and only if $\mathfrak{A}^{\prime}$ pp-constructs $\mathfrak{K}_{3}$.
  2. (2)
    $\mathcal{J}$ and $\mathcal{J}^{\prime}$ are (Datalog-)reductions from $\mathrm{CSP}(\mathfrak{A})$ to $\mathrm{CSP}(\mathfrak{A}^{\prime})$ and vice versa.
  3. (3)
    If $\mathfrak{A}$ is a model-complete core, then $\mathfrak{A}^{\prime}$ is a model-complete core as well.
  4. (4)
    If $\mathfrak{B}$ is Ramsey, then $\mathfrak{B}^{\prime}$ is Ramsey as well.
  5. (5)
    Every finite cheese for $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ pp-constructs $\mathfrak{K}_{3}$; hence, $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ is not finitely tractable.
  6. (6)
    $\mathfrak{A}^{\prime}$ is a cheese for $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$; hence, $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ reduces to $\mathrm{CSP}(\mathfrak{A})$ via $\mathcal{J}^{\prime}$.
  7. (7)
    Conversely, $\mathcal{J}$ is correct as a reduction from $\mathrm{CSP}(\mathfrak{A})$ to $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ when limited to instances $\mathfrak{X}$ of $\mathrm{CSP}(\mathfrak{A})$ with the property that $\mathfrak{X}$ homomorphically maps to $\mathfrak{A}$ if and only if $\mathfrak{X}$ homomorphically maps to a substructure of $\mathfrak{A}$ of domain size $\leq n$.
The last item in Theorem 2.6 is not necessary for obtaining non-finitely tractable PCSPs from Conjecture 1.4, but provides an additional link that makes it possible to transfer logical inexpressiblity results from infinite-domain CSPs to PCSPs. For example, the PCSPs provided by the theorem from $\mathrm{CSP}(\mathbb{Q};X)$, where
$X=\{(x,y,z)\mid x=y<z\text{ or }y=z<x\text{ or }z=x<y\}$,
will not be solvable in Fixed-Point logic with Counting (FPC), because $\mathrm{CSP}(\mathbb{Q};X)$ is not [33] and the instances witnessing the inexpressiblity in FPC homomorphically map to $(\mathbb{Q};X)$ if and only if they homomorphically map to a fixed finite substructure of it (cf. [57, Proposition 37]). In contrast, $\mathrm{CSP}(\mathbb{Q};\{(x,y,z)\mid x>y\text{ or }x>z\})$ is not solvable in Datalog [23], but it can be deduced from the results in [57] (Theorem 3) that the associated finite-domain PCSPs are.
As mentioned earlier, the first two examples of a tractable $\omega$-categorical cheese for a finite-domain PCSP that is not finitely tractable have recently been provided by Mottet [57, Propositions 35 and 36]. Our results compare to the results of that work as follows. First, Mottet’s finite-domain PCSPs [57, Theorem 2] are reducible to the associated $\omega$-categorical templates under gadget reductions, which are the specific form of Datalog-reductions stemming from pp-constructions. The main advantage of gadget reductions over general Datalog-reductions is that, in many settings, gadget reductions between (P)CSPs are captured by the transfer of height-1 identities between polymorphism clones (or more generally polymorphism minions); see [38]. As a trade-off for the generality of our result (Theorem 2.6), our Datalog-reductions are not given in terms of gadgets; in fact, one can formally prove that, in many cases, our finite-domain PCSPs do not admit a gadget reduction to the original $\omega$-categorical template (Example 4.6). Second, Mottet’s work [57] also contains a result (Theorem 1) showing that every CSP within the scope of the Bodirsky-Pinsker conjecture is polynomial-time equivalent to the PCSP of a half-infinite template, more precisely a template $(\mathfrak{S}_{1},\mathfrak{S}_{2})$ such that $\mathfrak{S}_{1}$ is in the scope of the Bodirsky-Pinsker conjecture and $\mathfrak{S}_{2}$ is finite. In contrast, our Theorem 2.6 provides non-finitely tractable finite-domain PCSP templates $(\mathfrak{S}_{1},\mathfrak{S}_{2})$ sandwiching a given structure $\mathfrak{A}$ from Conjecture 1.4 up to Datalog interreducibility. Since our PCSP templates $(\mathfrak{S}_{1},\mathfrak{S}_{2})$ are finite and not half-infinite, we cannot guarantee polynomial-time equivalence between $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ and $\mathrm{CSP}(\mathfrak{A})$. However, our construction still allows us to approximate $\mathrm{CSP}(\mathfrak{A})$ to an arbitrary degree in the sense of item 7 in Theorem 2.6.
3. Basic definitions
The set $\{1,\dots,n\}$ is denoted by $[n]$, and we use the bar notation $\bar{t}$ for tuples. For a function $f\colon A^{n}\rightarrow B$, the component-wise action of $f$ on $k$-tuples $(x_{1,1},\dots,x_{1,k}),$ $\dots$,… , $(x_{n,1},\dots,x_{n,k})\in A^{k}$ is defined as the $k$-tuple
$f\big{(}(x_{1,1},\dots,x_{1,k}),\dots,(x_{n,1},\dots,x_{n,k})\big{)}\coloneqq\big{(}f(x_{1,1},\dots,x_{n,1}),\dots,f(x_{1,k},\dots,x_{n,k})\big{)}.$
3.1. Relational structures.
A (relational) signature $\tau$ is a set of relation symbols, each $R\in\tau$ with an associated natural number called arity. A (relational) $\tau$-structure $\mathfrak{A}$ consists of a set $A$ (the domain) together with the relations $R^{\mathfrak{A}}\subseteq A^{k}$ for each $R\in\tau$ with arity $k$. An expansion of $\mathfrak{A}$ is a $\sigma$-structure $\mathfrak{B}$ with $A=B$ such that $\tau\subseteq\sigma$ and $R^{\mathfrak{B}}=R^{\mathfrak{A}}$ for each relation symbol $R\in\tau$. Conversely, we call $\mathfrak{A}$ a reduct of $\mathfrak{B}$ and denote it by $\mathfrak{B}|^{\tau}$. For a positive integer $d$ and a $\tau$-structure $\mathfrak{A}$, the $d$-th power of $\mathfrak{A}$ is the $\tau$-structure $\mathfrak{A}^{d}$ with domain $A^{d}$ and relations
$R^{\mathfrak{A}^{d}}=\{\big{(}(a_{1,1},\dots,a_{1,d}),\dots,(a_{k,1},\dots,a_{k,d})\big{)}\in(A^{d})^{k}\mid(a_{1,1},\dots,a_{k,1}),\dots,(a_{1,d},\dots,a_{k,d})\in R^{\mathfrak{A}}\}$
for each $R\in\tau$ of arity $k$. The substructure of a $\tau$-structure $\mathfrak{A}$ on a subset $B\subseteq A$ is the $\tau$-structure $\mathfrak{B}$ with domain $B$ and relations $R^{\mathfrak{B}}=R^{\mathfrak{A}}\cap B^{k}$ for every $R\in\tau$ of arity $k$. The factor of a $\tau$-structure $\mathfrak{A}$ through an equivalence relation $E\subseteq A^{2}$ is the $\tau$-structure $\mathfrak{A}/{E}$ with domain $A/E$ and relations $R^{\mathfrak{A}/{E}}=q_{E}(R^{\mathfrak{A}})$, where $q_{E}$ denotes the factor map $x\mapsto[x]_{E}$.
A homomorphism $h\colon\mathfrak{A}\rightarrow\mathfrak{B}$ for $\tau$-structures $\mathfrak{A}$ and $\mathfrak{B}$ is a mapping $h\colon A\rightarrow B$ that preserves each $\tau$-relation, i.e., if $\bar{t}\in R^{\mathfrak{A}}$ for some relation symbol $R\in\tau$, then $h(\bar{t})\in R^{\mathfrak{B}}$. We write $\mathfrak{A}\rightarrow\mathfrak{B}$ if $\mathfrak{A}$ maps homomorphically to $\mathfrak{B}$; formally, $\mathrm{CSP}(\mathfrak{A})=\{\mathfrak{J}\text{ finite}\mid\mathfrak{J}\rightarrow\mathfrak{A}\}$. If $\mathfrak{A}\rightarrow\mathfrak{B}$ and $\mathfrak{B}\rightarrow\mathfrak{A}$, then we say that $\mathfrak{A}$ and $\mathfrak{B}$ are homomorphically equivalent. An endomorphism of $\mathfrak{A}$ is a homomorphism from $\mathfrak{A}$ to itself; we denote by $\mathrm{End}(\mathfrak{A})$ the set (monoid) of all endomorphisms of $\mathfrak{A}$. Clearly, if $\mathfrak{A}$ has a constant endomorphism, then its CSP is trivial, and we call $\mathfrak{A}$ trivial itself; otherwise we call $\mathfrak{A}$ non-trivial. As mentioned in the introduction, a polymorphism of $\mathfrak{A}$ is a homomorphism from $\mathfrak{A}^{k}$ into $\mathfrak{A}$ for some $k\in\mathbb{N}$.
An embedding is an injective homomorphism $h\colon\mathfrak{A}\rightarrow\mathfrak{B}$ that additionally satisfies the following condition: for every $k$-ary relation symbol $R\in\tau$ and $\bar{t}\in A^{k}$ we have $h(\bar{t})\in R^{\mathfrak{B}}$ only if $\bar{t}\in R^{\mathfrak{A}}.$ An isomorphism is a surjective embedding. Two structures $\mathfrak{A}$ and $\mathfrak{B}$ are isomorphic if there exists an isomorphism from $\mathfrak{A}$ to $\mathfrak{B}$. An automorphism is an isomorphism from $\mathfrak{A}$ to $\mathfrak{A}$; we denote by $\mathrm{Aut}(\mathfrak{A})$ the set (group) of all automorphisms of $\mathfrak{A}$. The orbit of a tuple $\bar{t}\in A^{k}$ in $\mathfrak{A}$ is the set $\{g(\bar{t})\mid g\in\mathrm{Aut}(\mathfrak{A})\}$. Any tuples belonging to the same orbit satisfy precisely the same first-order formulas over $\mathfrak{A}$. The orbit equivalence relation provides a natural way to factorise relational structures; in the present paper, we will only need the following specific case. Given a relational structure $\mathfrak{A}$ and a subgroup $G\subseteq\mathrm{Aut}(\mathfrak{A})$, we denote by $\mathfrak{A}_{\!{}_{\vphantom{/}\smash{/G}}}$ the structure $\mathfrak{A}/E$ for
$E\coloneqq\{(x,y)\in A^{2}\mid\exists g\in G\colon g(x)=y\}.$
3.2. Universal algebra.
An (equational) condition is a set of identities, i.e. formal expressions of the form $s\approx t$, where $s$ and $t$ are terms over a common set of function symbols. An equational condition is height-1 if it contains neither nested terms nor terms consisting of a single variable. We say that $\mathrm{Pol}(\mathfrak{A})$ (or some set of operations on $A$) satisfies an equational condition $\Sigma$ if the function symbols can be interpreted as elements of $\mathrm{Pol}(\mathfrak{A})$ so that, for each identity $s\approx t$ in $\Sigma$, the equality $s=t$ holds for any evaluation of variables in $A$. An example is the cyclic identity of arity $n\geq 2$ given by $f(x_{1},\dots,x_{n})\approx f(x_{2},\dots,x_{n},x_{1})$, for a symbol $f$ of arity $n\geq 2$. We say that an operation is cyclic if it satisfies the cyclic identity.
Theorem 3.1 ([8]).
Let $\mathfrak{A}$ be a finite relational structure. If $\mathfrak{A}$ does not pp-construct $\mathfrak{K}_{3}$, then $\mathrm{Pol}(\mathfrak{A})$ contains a cyclic operation.
A pseudo-version of a height-1 identity $s\approx t$ is of the form $\alpha\circ s\approx\beta\circ t$ for fresh unary function symbols $\alpha$ and $\beta$. An example is the pseudo-Siggers identity from Conjecture 1.4. This definition naturally extends to height-1 conditions.
4. Simplifying the Bodirsky-Pinsker conjecture
4.1. Injectivity and Datalog reductions
In the context of Conjecture 1.4, the concepts of CSP-injectivity and having no algebraicity are closely related. On the one hand, if $\mathfrak{B}$ is a homogeneous structure with no algebraicity and its relations only contain tuples with pairwise distinct entries, then it is CSP-injective [15, Lemma 4.3.6]. On the other hand, we can prove the following.
Proposition 4.1.
CSP-injective $\omega$-categorical model-complete cores do not have algebraicity.
We shall now see that CSP-injectivity and Pol-injectivity cannot hold simultaneously in $\omega$-categorical structures; yet, we shall observe immediately thereafter that CSP-injectivity can be traded against Pol-injectivity using the relation $I_{4}$ from Example 2.1.
Proposition 4.2.
CSP-injective $\omega$-categorical structures are not Pol-injective.
Proposition 4.3.
The expansion $\mathfrak{A}_{I_{4}}$ of any CSP-injective structure $\mathfrak{A}$ by the relation $I_{4}$ is Pol-injective, and $\mathrm{CSP}(\mathfrak{A})$ and $\mathrm{CSP}(\mathfrak{A}_{I_{4}})$ are Datalog-interreducible.
Datalog reductions between CSPs are typically specified using factorisation-free Datalog-interpretations with parameters [1, 38], which are a particular case of logical interpretations with parameters. In the present paper, we only need a very basic understanding of Datalog-reductions, which can be easily explained on an informal level. Datalog is defined by adding formation rules to the existential positive fragment of first-order logic whose semantics is defined with inflationary fixed-points of definable operators. Every existential positive first-order formula is a Datalog formula and, if $\phi(\bar{x})$ is a Datalog formula over some relational signature $\tau\cup\{R\}$ with $R\notin\tau$, then $[\mathsf{ifp}_{R}\,\phi](\bar{x})$ is a Datalog formula over the signature $\tau$ whose semantics is given as follows. For a $\tau$-structure $\mathfrak{A}$ and a tuple $\bar{a}$ over $A$ matching the arity of $\bar{x}$, say $k$, we have $\mathfrak{A}\models[\mathsf{ifp}_{R}\,\phi](\bar{a})$ if and only if $\bar{a}$ is contained in the inflationary fixed-point of the operator $F^{\mathfrak{A}}_{\phi,R}(X)\coloneqq\{\bar{x}\in A^{k}\mid\mathfrak{A}_{X}\models\phi(\bar{x})\}$, i.e., the limit of the sequence $X_{0}\coloneqq\emptyset$ and $\smash{X_{i+1}\coloneqq X_{i}\cup F^{\mathfrak{A}}_{\phi,R}(X_{i})}$. For example, the Datalog formula
$\phi_{<}^{tc}(x,z)\coloneqq[\mathsf{ifp}_{T}\,(x<z)\vee(\exists y\ldotp T(x,y)\wedge T(y,z))](x,z)$
computes the transitive closure of $<$.
For a relational $\tau$-structure $\mathfrak{A}$, we say that $\mathrm{CSP}(\mathfrak{A})$ is solvable in Datalog if there exists a Datalog sentence defining the class of all finite $\tau$-structures which do not homomorphically map to $\mathfrak{A}$. For example, $\mathrm{CSP}(\mathbb{Q};<)$ is solvable in Datalog using the sentence $\exists u\ldotp\phi_{<}^{tc}(u,u)$. It is not hard to see that every Datalog formula is specified by a finite set of rules of the form $R(\bar{x})\Leftarrow R_{1}(\bar{x}_{1})\wedge\cdots\wedge R_{m}(\bar{x}_{m})$ where $R$ is a fixed-point variable. In the case of the above Datalog formula computing the transitive closure of $<$, for example, the rules are $T(x,z)\Leftarrow(x<z)$ and $T(x,z)\Leftarrow T(x,y)\wedge T(y,z)$. When it comes to Datalog-reductions between CSPs, we only need the following intuitive understanding of this concept.
A Datalog-interpretation is a mapping $\mathcal{I}$ from finite $\sigma$-structures to finite $\tau$-structures such that for every finite $\sigma$-structure $\mathfrak{J}$ the $\tau$-structure $\mathcal{I}(\mathfrak{J})$ can be defined from $\mathfrak{J}$ or a fixed finite power of it using Datalog formulas [1, Definition 1]. For a $\tau$-structure $\mathfrak{A}$ and $\sigma$-structure $\mathfrak{A^{\prime}}$, we say that $\mathrm{CSP}(\mathfrak{A^{\prime}})$ Datalog-reduces to $\mathrm{CSP}(\mathfrak{A})$ if there exists a Datalog interpretation $\mathcal{I}$ so that $\mathfrak{J}\rightarrow\mathfrak{A^{\prime}}$ if and only if $\mathcal{I}(\mathfrak{J})\rightarrow\mathfrak{A}$. Datalog reductions compose because Datalog interpretations do, i.e., if $\mathrm{CSP}(\mathfrak{A}^{\prime\prime})$ Datalog-reduces to $\mathrm{CSP}(\mathfrak{A}^{\prime})$ and $\mathrm{CSP}(\mathfrak{A}^{\prime})$ Datalog-reduces to $\mathrm{CSP}(\mathfrak{A})$, then $\mathrm{CSP}(\mathfrak{A}^{\prime\prime})$ Datalog-reduces to $\mathrm{CSP}(\mathfrak{A})$.
4.2. Wreath products
Let $\mathfrak{A}$ and $\mathfrak{B}$ be structures over disjoint relational signatures $\tau$ and $\sigma$. Moreover, let $E$ be a fresh binary symbol. The wreath product $\mathfrak{A}\wr\mathfrak{B}$ is the structure over the signature $\tau\cup\{E\}\cup\sigma$ with domain $A\times B$ whose relations are defined as follows. First,
$E^{\mathfrak{A}\wr\mathfrak{B}}\coloneqq\{((x_{1},y_{1}),(x_{2},y_{2}))\in(A\times B)^{2}\mid y_{1}=y_{2}\}.$
Then, for every $R\in\tau$ of arity $k$:
$R^{\mathfrak{A}\wr\mathfrak{B}}\coloneqq\{((x_{1},y_{1}),\dots,(x_{k},y_{k}))\in(A\times B)^{k}\mid(x_{1},\dots,x_{k})\in R^{\mathfrak{A}}\text{ and }y_{1}=\cdots=y_{n}\},$
and, for every $S\in\sigma$ of arity $m$:
$S^{\mathfrak{A}\wr\mathfrak{B}}\coloneqq\{((x_{1},y_{1}),\dots,(x_{m},y_{m}))\in(A\times B)^{m}\mid(y_{1},\dots,y_{m})\in S^{\mathfrak{B}}\}.$
Below we give several crucial properties of wreath products, but first we start with an illustratory example on which all of these properties can be readily verified.
Example 4.4.
Consider the wreath product $(\mathbb{Q};<_{1})\wr(\mathbb{Q};<_{2})$ of two isomorphic copies of $(\mathbb{Q};<)$ with disjoint signatures. The union $<_{\text{conv}}$ of $<_{1}$ and $<_{2}$ in this structure, together with $E$, is the random convexly ordered equivalence relation from Corollary 6.8 in [52].
Given groups $G$ and $H$ acting on sets $A$ and $B$, respectively, their wreath product $G\wr H$ is given by $G^{B}\times H$ acting on $A\times B$ via $\left((g_{b})_{b\in B},h)\right)(a,b)\coloneqq(g_{h(b)}(a),h(b)).$ The wreath product $G\wr H$ again forms a group, with the neutral element $((e_{G})_{b\in B},e_{H})$ and the group operation
$((g_{b})_{b\in B},h))\cdot((g^{\prime}_{b})_{b\in B},h^{\prime}))\coloneqq((g_{b}\cdot g^{\prime}_{h^{-1}(b)})_{b\in B},h\cdot h^{\prime}))$.
Note that the permutation group on $A\times B$ induced by the component-wise action of the Cartesian product $G\times H$ is contained by that induced by the action of $G\wr H$ thereon: the permutations induced stemming from the former are represented in $G\wr H$ by all elements of the form $((g)_{b\in B},h)$.
Proposition 4.5 (Basic properties of wreath products).
Let $\mathfrak{A}$ and $\mathfrak{B}$ be structures over disjoint finite relational signatures $\tau$ and $\sigma$. Then:
  1. (1)
    $\mathrm{Aut}(\mathfrak{A}\wr\mathfrak{B})=\mathrm{Aut}(\mathfrak{A})\wr\mathrm{Aut}(\mathfrak{B})$.
  2. (2)
    If $\mathfrak{A}$ has no algebraicity, then $\mathfrak{A}\wr\mathfrak{B}$ has no algebraicity.
  3. (3)
    If $\mathfrak{A}$ and $\mathfrak{B}$ are $\omega$-categorical, then $\mathfrak{A}\wr\mathfrak{B}$ is $\omega$-categorical.
  4. (4)
    If $\mathfrak{A}$ and $\mathfrak{B}$ are homogeneous, then $\mathfrak{A}\wr\mathfrak{B}$ is homogeneous.
  5. (5)
    If $\mathfrak{A}$ and $\mathfrak{B}$ are finitely bounded, then $\mathfrak{A}\wr\mathfrak{B}$ is finitely bounded.
  6. (6)
    If $\mathfrak{A}$ and $\mathfrak{B}$ are homogeneous Ramsey, then $\mathfrak{A}\wr\mathfrak{B}$ is homogeneous Ramsey.
Item 6 in Proposition 4.5 follows from a famous theorem of Kechris, Pestov, and Todorčević [52] stating that a homogeneous structure $\mathfrak{B}$ with a countable relational signature is Ramsey if and only if its automorphism group $\mathrm{Aut}(\mathfrak{B})$ is extremely amenable, and hence the Ramsey property only depends on $\mathrm{Aut}(\mathfrak{B})$ viewed as an abstract topological group. Kechris, Pestov, and Todorčević observed that extreme amenability is preserved under taking wreath products; although they do not state this result explicitly, they do provide an example as a proof of concept: the automorphism group of the random convexly ordered equivalence relation (Example 4.4). Scow [64, Theorem 5.13] does state this result, and essentially provides a full proof, modulo Theorem 4.7 in [52].
4.3. Removing algebraicity
Proof sketch of Theorem 2.2 and Theorem 2.4.
Let $\tau$ and $\sigma$ be the signatures of $\mathfrak{A}$ and $\mathfrak{B}$, respectively. We first define the auxiliary structure $\mathfrak{B}^{\prime}\coloneqq(\mathbb{Q};<)\wr\mathfrak{B}$. As shown in Examples 1.3 and 2.1, the structure $(\mathbb{Q};<)$ is finitely bounded, homogeneous, and has no algebraicity. It then follows immediately from Proposition 4.5(2) that $\mathfrak{B}^{\prime}$ has no algebraicity. The expansion $\mathfrak{B}^{\prime}_{\neq}$ of $\mathfrak{B}^{\prime}$ by the binary inequality has no algebraicity because taking expansions by first-order definable relations does not change the automorphism group; the same is true for the structure $\mathfrak{B}^{\prime}_{\neq,I_{4}}$. Note that removing all symbols in $\sigma\setminus\tau$ as well as the symbol $<$ from $\mathfrak{B}^{\prime}_{\neq}=\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq}$ yields $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$.
We claim that $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ is CSP-injective. To this end, let $\mathfrak{J}$ be an arbitrary finite structure for which there exists a homomorphism $h\colon\mathfrak{J}\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. We define $\mathfrak{J}_{\pi}$ as the $(\tau\cup\{E\})$-structure with domain $J$ such that, for every $R\in\tau\cup\{E\}$, we have $\bar{t}\in R^{\mathfrak{J}_{\pi}}$ if and only if $h(\bar{t})\in R^{(\mathbb{Q}\wr\mathfrak{A})_{\neq}}$. It is easy to see that $\mathfrak{J}_{\pi}$ embeds into the $(\tau\cup\{E\})$-reduct of $\mathfrak{B}^{\prime}$ (note that if multiple elements in $J$ are mapped to $(q,a)$ for some $q\in\mathbb{Q}$ and $a\in A$, we can always map them to different elements in ×$\mathbb{Q}\times\{a\}$). Hence, $\mathfrak{J}_{\pi}$ expanded by the binary inequality predicate embeds into $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ and admits an injective homomorphism from $\mathfrak{J}$. We conclude that $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ is CSP-injective.
Next, we informally describe two Datalog-reductions $\mathcal{I}^{\ast}$ and $\mathcal{I}$: from $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$ to $\mathrm{CSP}(\mathfrak{A})$ and back, starting with the former. Let $\mathfrak{J}$ be an instance of $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$. Denote by $\sim$ the equivalence closure of $E^{\mathfrak{J}}$, and let $q_{\sim}\colon\mathfrak{J}\rightarrow(\mathfrak{J}/{\sim}),x\mapsto[x]_{\sim}$ be the induced quotient map. We define $\mathcal{I}^{\ast}(\mathfrak{J})$ as the $\tau$-structure with domain $J$ and whose relations are the preimages of all $\tau$-relations in $\mathfrak{J}/{\sim}$ under $q_{\sim}$, except that we additionally add, for every $x\in J$ such that $(x,x)\in{\neq}^{\mathfrak{J}}$, the tuple $(x,\dots,x)$ to every $\tau$-relation. It is easy to see that $\mathcal{I}^{\ast}(\mathfrak{J})$ is definable in Datalog; we verify that $\mathcal{I}^{\ast}$ is a reduction from $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$ to $\mathrm{CSP}(\mathfrak{A})$. Clearly, if $\mathfrak{J}$ maps homomorphically to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$, then $\mathcal{I}^{\ast}(\mathfrak{J})$ maps homomorphically to $\mathfrak{A}$. Suppose that there exists a homomorphism $h\colon\mathcal{I}^{\ast}(\mathfrak{J})\rightarrow\mathfrak{A}$. Since by non-triviality $\mathfrak{A}$ does not have a constant endomorphism, there is no $x\in J$ such that $(x,x)\in{\neq}^{\mathfrak{J}}$. Denote by $\mathfrak{J}^{\prime}$ the structure obtained from $\mathfrak{J}$ by removing all tuples from ${\neq}^{\mathfrak{J}}$. Now, choosing representatives $j_{1},\dots,j_{k}$ for all classes of $\mathfrak{J}/{\sim}$, the function sending the entirety of $[j_{i}]_{\sim}$ to $(0,h(j_{i}))$ for all $i\leq k$ is a homomorphism from $\mathfrak{J}^{\prime}$ to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. By the CSP-injectivity of $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$, it can be replaced by an injective homomorphism, which is then also a homomorphism from $\mathfrak{J}$ to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. We continue with the reduction from $\mathrm{CSP}(\mathfrak{A})$ to $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$, which is trivial. Let $\mathfrak{J}$ be an instance of $\mathrm{CSP}(\mathfrak{A})$. We define $\mathcal{I}(\mathfrak{J})$ as the $(\tau\cup\{E,\neq\})$-expansion of $\mathfrak{J}$ by empty relations. By CSP-injectivity, there exists a homomorphism $h\colon\mathcal{I}(\mathfrak{J})\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ if and only if there exists an injective such homomorphism $i$. This is the case if and only if the relation $E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq}}$ restricted to $i(J)$ describes the kernel of a homomorphism from $\mathfrak{J}$ to $\mathfrak{A}$. Therefore, $\mathfrak{J}\rightarrow\mathfrak{A}$ if and only if $\mathcal{I}(\mathfrak{J})\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$.
We now consider $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$. By Proposition 4.3, $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ is Pol-injective and $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}})$ and $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$ are Datalog-interreducible. That $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}})$ and $\mathrm{CSP}(\mathfrak{A}$) are Datalog-interreducible then follows from the fact that Datalog-reductions can be composed.
Finally, we sketch how to verify the properties in items 123, and 4. We only explicitly cover the case where $I_{4}$ is present (Theorem 2.4), since the arguments without $I_{4}$ are a subset.
Regarding item 1, the first part ($\omega$-categoricity) follows directly from Proposition 4.5(3) applied to $\mathfrak{B}^{\prime}$ and since we take an expansion by relations which are preserved by all bijections, and in particular all automorphisms. The second part of item 1 (model-complete cores) is immediate once one observes that the endomorphism monoid of a wreath product admits a similar description as its automorphism group, cf. Proposition 4.5(1).
Regarding item 2 (homogeneity and the Ramsey property), one first observes that, since $\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq,I_{4}}$ is an expansion of $\mathfrak{B}^{\prime}$ by two relations that are preserved by all embeddings, we can ignore these relations and only prove the statement for $\mathfrak{B}^{\prime}$. Then homogeneity follows directly from Proposition 4.5(4) because $(\mathbb{Q};<)$ is homogeneous. In the second part, the backward direction follows directly from Proposition 4.5(6) because $(\mathbb{Q};<)$ is homogeneous Ramsey [52]. The forward direction can be proved by hand, by instantiating the Ramsey property for the finite substructures of $\mathfrak{B}^{\prime}$ where $E$ interprets as the diagonal relation.
Regarding item 3 (finite boundedness), this is a consequence of Proposition 4.5(5) (which immediately yields finite boundedness of $\mathfrak{B}^{\prime}$) and the fact that we take an expansion by relations definable by a Boolean combination of equality atoms. More precisely, recall that up to isomorphism, the finite substructures of a finitely bounded structure are precisely the finite models of a universal sentence. By adding new clauses defining the relations $\neq$ and $I_{4}$ in terms of equalities to the universal sentence describing the finite substructures of $\mathfrak{B}^{\prime}$, we obtain a universal sentence describing the finite substructures of $\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq,I_{4}}$.
Regarding item 4 (pp-construction of $\mathfrak{K}_{3}$), we use a characterization via the pseudo-Siggers identity from [7, Theorems 1.3 and 3.4] and the fact that the structure $(\mathbb{Q};\neq,I_{4})$ is an $\omega$-categorical model-complete core that does not pp-construct $\mathfrak{K}_{3}$ [15, Theorems 12.0.1, 12.7.3, 12.9.2, and Corollary 6.4.4]. One direction is simple (and does not use the assumption on the orbit growth), because $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ clearly pp-constructs $\mathfrak{A}$; the pp-construction is given by $\mathfrak{A}=(\smash{(\mathbb{Q}\wr\mathfrak{A})_{\neq}/E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}})|^{\tau}}$ (since taking factors and reducts are pp-constructions [9]). Thus, if $\mathfrak{A}$ pp-constructs $\mathfrak{K}_{3}$, then so does $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$. For the other direction, one assumes that $\mathfrak{A}$ does not pp-construct $\mathfrak{K}_{3}$, in which case Theorem 1.3 in [7] yields polymorphisms in its model-complete core satisfying the pseudo-Siggers identity. These are then composed with polymorphisms of $(\mathbb{Q};\neq,I_{4})$ witnessing the pseudo-Siggers identity, which ultimately yields the satisfaction of the pseudo-Siggers identity by polymorphisms of the model-complete core of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$. The proof of Proposition 4.5(3) shows that also the structure $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ has less than double exponential orbit growth. It then follows from Theorem 3.4 in [7] (whose item (ii) is equivalent to the pp-construction of $\mathfrak{K}_{3}$ by Theorem 1.8 in [9]) that $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ does not pp-construct $\mathfrak{K}_{3}$. ∎
We conclude this section with an example.
Example 4.6.
Consider the two structures
$\mathfrak{A}_{1}\coloneqq(\{0,1\};\{1\},\{(x,y,z)\in\{0,1\}^{3}\mid x+y+z=0\bmod 2\}),$
$\mathfrak{A}_{2}\coloneqq(\mathbb{Q};<,\{(x,y,z)\in\mathbb{Q}^{3}\mid x\geq y\text{ or }x\geq z\}).$
The former is a finitely bounded homogeneous model-complete core with algebraicity (due to it being finite) and the latter is a finitely bounded homogeneous model-complete core without algebraicity. Note that $\mathfrak{A}_{1}$ is preserved by the symmetric operation $(x_{1},\dots,x_{n})\mapsto x_{1}+\cdots+x_{n}\bmod 2$ for every odd $n\geq 2$ and $\mathfrak{A}_{2}$ is preserved by the symmetric operation $(x_{1},\dots,x_{n})\mapsto\min\{x_{1},\dots,x_{n}\}$ for every $n\geq 2$. Thus, in both cases the polymorphism clone satisfies the cyclic identity of arity $n$ for some $n\geq 2$. Since pp-constructions preserve the satisfaction of height-1 identities in polymorphism clones [9] and no cyclic operation $f$ of arity $n\geq 2$ preserves $\neq$ over an infinite set, $\mathfrak{A}_{i}$ does not pp-construct $(\mathbb{Q}\wr\mathfrak{A}_{i})_{\neq}$ for both $i\in[2]$. Hence, our Datalog-reduction in Theorem 2.2 is not subsumed by gadget reductions (pp-constructions).
5. $\omega$-categorical cheeses for PCSPs
5.1. PCSP polymorphisms and full powers
We extend the notion of a polymorphism to PCSP templates $(\mathfrak{S}_{1},\mathfrak{S}_{2})$ in the usual way: a polymorphism of $(\mathfrak{S}_{1},\mathfrak{S}_{2})$ is a homomorphism from a finite power of $\mathfrak{S}_{1}$ to $\mathfrak{S}_{2}$; we denote by $\mathrm{Pol}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ the set (minion) of all such polymorphisms. Also the satisfaction of height-1 conditions can be generalized to the PCSP setting in a straightforward manner. In the case of nested identities, one has to be more careful because the composition of polymorphisms between different structures is not defined.We say that a height-1 condition $\Sigma$ is satisfied in $\mathrm{Pol}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ modulo a set $H$ of unary operations on $S_{2}$ if the function symbols in $\Sigma$ can be interpreted as elements of $\mathrm{Pol}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ such that for each identity $s\approx t$ in $\Sigma$ there exist $\alpha,\beta\in H$ such that $\alpha\circ s=\beta\circ t$ holds for any evaluation of variables in $S_{1}$.
Our basic tool for creating $\omega$-categorical cheeses for PCSPs that are not finitely tractable are full powers (see, e.g., Bodirsky [15, Section 3.5]). Roughly speaking, a full power ${\mathfrak{A}}^{[d]}$ is a higher-dimensional representation of $\mathfrak{A}$ that can be pp-constructed from it and vice versa; its central feature is that its polymorphisms are the polymorphisms of $\mathfrak{A}$ acting on $d$-tuples, thus allowing to factor it by the orbit-equivalence of $\mathrm{Aut}(\mathfrak{A})$ on $d$-tuples (rather than on elements). In other words, the orbits of $d$-tuples of $\mathfrak{A}$ are represented by orbits of $1$-tuples of ${\mathfrak{A}}^{[d]}$, which is crucial in applications of the sandwich method. Let $\mathfrak{A}$ be a relational structure with a signature $\tau$ and let $d\in\mathbb{N}$ be arbitrary. The $d$-th full power of $\mathfrak{A}$, denoted ${\mathfrak{A}}^{[d]}$, is the structure with domain $A^{d}$ and the following relations for every $k\in[d]$:
  • for every $R\in\tau$ of arity $k$ and every injection $\iota\colon[k]\rightarrow[d]$, the unary relation
    $R^{{\mathfrak{A}}^{[d]}}_{\iota}\coloneqq\{(a_{1}\dots,a_{d})\in A^{d}\mid(a_{\iota(1)},\dots,a_{\iota(k)})\in R^{\mathfrak{A}}\};$
  • for every $R\in\tau$ of arity $k$ and every function $\iota\colon[k]\rightarrow[d]$, the $k$-ary relation
    $\widehat{R}^{{\mathfrak{A}}^{[d]}}_{\iota}\coloneqq\{((a^{1}_{1},\dots,a^{1}_{d}),\dots,(a^{k}_{1},\dots,a^{k}_{d}))\in(A^{d})^{k}\mid(a^{1}_{\iota(1)},\dots,a^{k}_{\iota(k)})\in R^{\mathfrak{A}}\};$
  • for all functions $\iota,\iota^{\prime}\colon[k]\rightarrow[d]$, the binary compatibility relation
    $S^{{\mathfrak{A}}^{[d]}}_{\iota,\iota^{\prime}}\coloneqq\{((a^{1}_{1}\dots,a^{1}_{d}),(a^{2}_{1},\dots,a^{2}_{d}))\in(A^{d})^{2}\mid\forall i\in[k]\colon a^{1}_{\iota(i)}=a^{2}_{\iota^{\prime}(i)}\}.$
Below we give several important properties of full powers.
Proposition 5.1 (Basic properties of full powers).
Let $\mathfrak{A}$ be a relational structure over a finite signature $\tau$ and let its relations be of arity $\leq d$. Then:
  1. (1)
    $\mathrm{Pol}({\mathfrak{A}}^{[d]})$ consists of the component-wise actions of polymorphisms of $\mathfrak{A}$ on $d$-tuples.
  2. (2)
    ${\mathfrak{A}}^{[d]}$ is (factorisation-free) pp-interpretable from $\mathfrak{A}$ and vice versa.
  3. (3)
    If $\mathfrak{A}$ is $\omega$-categorical, then ${\mathfrak{A}}^{[d]}$ is $\omega$-categorical.
  4. (4)
    If $\mathfrak{A}$ is homogeneous, then ${\mathfrak{A}}^{[d]}$ is homogeneous.
  5. (5)
    If $\mathfrak{A}$ is a model-complete core, then ${\mathfrak{A}}^{[d]}$ is a model-complete core.
  6. (6)
    If $\mathfrak{A}$ is finitely bounded, then ${\mathfrak{A}}^{[d]}$ is finitely bounded.
  7. (7)
    If $\mathfrak{A}$ is homogeneous Ramsey, then ${\mathfrak{A}}^{[d]}$ is homogeneous Ramsey.
  8. (8)
    If $\mathfrak{A}$ is a reduct of $\mathfrak{B}$, then ${\mathfrak{A}}^{[d]}$ is a reduct of ${\mathfrak{B}}^{[d]}$.
5.2. Our sandwich recipe
Proposition 5.2.
Let $\mathfrak{A}$ be a reduct of a linearly ordered finitely bounded homogeneous structure $\mathfrak{B}$ such that one of the relations of $\mathfrak{A}$ interprets as $\neq$. Let $d\in\mathbb{N}$ be such that the bounds of $\mathfrak{B}$ have size $\leq d$ and its relations are of arity $\leq d-1$. Then, for every finite substructure $\mathfrak{S}$ of $\mathfrak{A}$ with $|S|\geq 3$:
  1. (1)
    ${\mathfrak{A}}^{[d]}$ is a cheese for $\mathrm{PCSP}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$;
  2. (2)
    Every finite cheese for $\mathrm{PCSP}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ pp-constructs $\mathfrak{K}_{3}$.
Proof sketch.
Clearly, for every finite substructure $\mathfrak{S}$ of $\mathfrak{A}$, we have $\smash{{\mathfrak{S}}^{[d]}\rightarrow{\mathfrak{A}}^{[d]}\rightarrow{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}}$. Now suppose that there is a finite cheese $\mathfrak{D}$ of $({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ that does not pp-construct $\mathfrak{K}_{3}$. Then, by Theorem 3.1, there exists a cyclic operation $f$ in $\mathrm{Pol}(\mathfrak{D})$ of some arity $n$. Take any homomorphisms $g\colon{\mathfrak{S}}^{[d]}\rightarrow\mathfrak{D}$ and $h\colon\mathfrak{D}\rightarrow{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$. Then, by setting $f^{\prime}(x_{1},\dots,x_{n})\coloneqq h\circ f(g(x_{1}),\dots,g(x_{n})),$ we get a cyclic $f^{\prime}\in\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ of arity $n$. One can show that every polymorphism from ${\mathfrak{S}}^{[d]}$ to ${\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$ can be lifted to a polymorphism from $\mathfrak{S}$ to $\mathfrak{A}$ and that any height-1 identity that is satisfied in $\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ is also satisfied in $\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ modulo $\mathrm{Aut}(\mathfrak{B})$. Hence, there is some $n$-ary $g\in\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ that is pseudo-cyclic modulo $\mathrm{Aut}(\mathfrak{B})$. One easily sees that any operation that is pseudo-cyclic modulo some order-preserving automorphism group must be cyclic. This means that we have successfully lifted the cyclic identity from $\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ to $\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$. But any cyclic operation on a set of size at least 3 does not preserve $\neq$, a contradiction. ∎
The following statement is folklore; see e.g. [15, Section 2.3.6].
Proposition 5.3.
[…]
Proof sketch (Theorem 2.6).
[…]
We select a finite substructure $\mathfrak{S}$ of $\mathfrak{A}$ with $|S|\geq 3$ large enough so that every finite structure with a homomorphism to a substructure of $\mathfrak{A}$ of size $\leq n$ also has a homomorphism to $\mathfrak{S}$; let $\widehat{\mathfrak{S}}$ be the substructure of $\widehat{\mathfrak{A}}$ induced by $\{0\}\times S$. We claim that $\mathfrak{A}^{\prime}\coloneqq{\widehat{\mathfrak{A}}}^{[d]}$, $\mathfrak{B}^{\prime}\coloneqq{\widehat{\mathfrak{B}}}^{[d]}$, $\mathfrak{S}_{1}\coloneqq{\widehat{\mathfrak{S}}}^{[d]}$, and $\mathfrak{S}_{2}\coloneqq{\widehat{\mathfrak{A}}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\widehat{\mathfrak{B}})}}}$ for $d$ as in Proposition 5.2 witness items 17 of the theorem. Items 1 to 6 follow from Theorem 2.2, Proposition 5.1, and Proposition 5.2. For item 7, note that if $\mathfrak{X}$ maps homomorphically to a substructure of $\mathfrak{A}$ of size $\leq n$, then $\mathfrak{X}\rightarrow\mathfrak{S}$, hence $\mathcal{I}(\mathfrak{X})\rightarrow\widehat{\mathfrak{S}}$ and hence $\mathcal{J}(\mathfrak{X})={\mathcal{I}(\mathfrak{X})}^{[d]}\rightarrow{\widehat{\mathfrak{S}}}^{[d]}=\mathfrak{S}_{1}$. Conversely, if $\mathfrak{X}\nrightarrow\mathfrak{A}$, then $\mathcal{I}(\mathfrak{X})\nrightarrow\widehat{\mathfrak{A}}$ by Theorem 2.2. One can show that if ${\mathcal{I}(\mathfrak{X})}^{[d]}\rightarrow\mathfrak{S}_{2}$, then also $\mathcal{I}(\mathfrak{X})\rightarrow\widehat{\mathfrak{A}}$, hence $\mathcal{J}(\mathfrak{X})\nrightarrow\mathfrak{S}_{2}$. ∎
6. Conclusion and outlook
6.1. Topology seems irrelevant
In the context of infinite-domain CSPs, the influence of topological properties of the polymorphism clones of $\omega$-categorical structures has received much attention, with various results claiming their relevance [27] or irrelevance [11]: while generally, whether or not an $\omega$-categorical structure $\mathfrak{A}$ pp-constructs $\mathfrak{K}_{3}$ (or pp-interprets another $\omega$-categorical structure $\mathfrak{A^{\prime}}$) does depend on such topological properties [9, 30, 46, 32], in many situations it does not (see, e.g., [41]). By moving from any $\omega$-categorical model-complete core template to one without algebraicity, Theorem 2.2 allows us to restrict Conjecture 1.4 to a class of topologically well-behaved structures. Namely, for those recent research suggests that often the algebraic structure determines the topological one; at least this is a fact for the endomorphism monoid, where the relevant topology (pointwise convergence) can then be defined from the algebraic structure of the monoid alone [62]. For $\omega$-categorical structures with algebraicity, on the other hand, various examples [42, 18, 62] show that this need not be the case. This also implies that there is no hope of lifting isomorphisms between endomorphism monoids or polymorphism clones of $\omega$-categorical structures to their counterparts constructed in Theorem 2.2.
6.2. Necessity of cheese preprocessing
In our proof of Theorem 2.6, we apply Proposition 5.2 to a structure which we previously “preprocessed” using Theorem 2.2 (removing algebraicity) and generic superpositions. It is natural to ask how strong Proposition 5.2 is on its own: can a result similar to Theorem 2.6 be proved directly using this proposition, perhaps under a more general structural assumption, e.g. that $\mathfrak{A}$ is a model-complete core? Here the answer seems to be negative. Consider the model-complete core structure
$\mathfrak{A}=(\mathbb{Q};<,\{(x,y,z)\in\mathbb{Q}^{3}\mid x\geq y\text{ or }x\geq z\})$
within the scope of Conjecture 1.4 ($\mathfrak{A}_{2}$ in Example 4.6); its CSP is polynomial-time tractable [22]. Let $\mathfrak{S}$ be any finite structure for which there exists a homomorphism $h\colon\mathfrak{S}\rightarrow{\mathfrak{A}}^{[d]}$, and let $\mathfrak{B}$ be an arbitrary $\omega$-categorical expansion of $\mathfrak{A}$. Let $A_{S}$ be the set of the elements in $A$ that appear as an entry of some element of $h(S)$, and let $\mathfrak{A}_{S}$ be the substructure of $\mathfrak{A}$ on $A_{S}$. Since $\smash{{\mathfrak{A}}^{[d]}\to{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}}$, the structure $\smash{{\mathfrak{A}}^{[d]}_{S}}$ is a finite cheese for $\smash{\mathrm{PCSP}(\mathfrak{S},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})}$.
We claim that $\mathrm{CSP}({\mathfrak{A}}^{[d]}_{S})$ is tractable, and hence this PCSP is finitely tractable. To see this, observe that the binary minimum operation $\min(x,y)$ is a polymorphism of $\mathfrak{A}$. This operation is conservative: $\min(x,y)\subseteq\{x,y\}$ holds for all $x,y$. Hence, its restriction to $A_{S}^{2}$ induces a polymorphism of $\mathfrak{A}_{S}$, which itself induces a polymorphism of $\smash{{\mathfrak{A}}^{[d]}_{S}}$ through its component-wise action. Since this operation is cyclic, $\smash{{\mathfrak{A}}^{[d]}_{S}}$ does not pp-construct $\mathfrak{K}_{3}$ [9, Theorem 1.4], and hence we are done by Theorem 1.1. This issue cannot be fixed simply by taking an expansion of $\mathfrak{A}$ by the inequality relation $\neq$ because $\mathrm{CSP}(\mathfrak{A};\neq)$ is NP-complete [22].
6.3. Algorithmic properties of our sandwiches
An interesting open question is whether the PCSP templates generated by Theorem 2.6 can be solved by some universal algorithm. Many known PCSP problems admitting infinite tractable cheeses can be solved by numeric relaxation algorithms such as BLP (the basic linear programming relaxation), AIP (the basic affine integer relaxation) or BLP+AIP (a combination of both) (see, e.g., [6, 34]). The proof of Proposition 5.2 shows that none of the PCSP templates produced by Theorem 2.6 admits a cyclic polymorphism. Since solvability of PCSPs by BLP is characterized by admitting totally symmetric polymorphisms of all arities [6, Theorem 7.9], which would also be cyclic, BLP does not solve any of those PCSPs. Regarding the combination BLP+AIP, we remark that replacing $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ and $\smash{\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq}}$ in the proof of Theorem 2.6 by $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ and $\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq,I_{4}}$ ensures that there is no cyclic and no 2-block symmetric operation in $\mathrm{Pol}(\mathfrak{S},(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}})$. In fact, we get the following much stronger statement, which follows directly from the fact that we take an expansion by $I_{4}$.
Proposition 6.1 (cf. proof of Lemma 7.5.1 in [15]).
Let $\Sigma$ be any finite height-1 condition that cannot be satisfied by any set of essentially injective operations from an at least 2-element set $S_{1}$ to a countably infinite set $S_{2}$. Then, for every at least 2-element substructure $\mathfrak{S}$ of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$, we have that $\mathrm{Pol}(\mathfrak{S},(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}})$ does not satisfy $\Sigma$. ∎
[…]
The lifting method we and Mottet [57] use provably cannot be extended to all height-1 conditions covered by Proposition 6.1. To see this, note that the Olšák identities
$f(y,x,x,x,y,y)\approx f(x,y,x,y,x,y)\approx f(x,x,y,y,y,x)$
cannot be satisfied by essentially injective operations but every finite-domain PCSP template without an Olšák polymorphism must be NP-hard [6, Corollary 6.3]. This means that the Olšák identities cannot possibly be prevented in our PCSPs without losing solvability in polynomial time (unless P=NP).
Appendix A Algebraicity is irrelevant (for infinite-domain CSPs)
A.1. Omitted definitions
We expand on some of the definitions appearing in the paper that were, for the sake of brevity, not presented in much detail. We also add the concept of a relational congruence, which will be useful in the the presentation of the proofs of Proposition 4.5, Lemma B.2, and Theorem 2.2, and state the theorem of Kechris, Pestov, and Todorčević referred to below Proposition 4.5, which is needed in the proof of Proposition 4.5.
Let $\mathfrak{A}$ be a structure over a finite relational signature containing $\tau$. The age of $\mathfrak{A}$, denoted by $\mathrm{age}(\mathfrak{A})$, is the class of all finite structures which embed into $\mathfrak{A}$. An equivalence relation $E$ over $A$ is a relational congruence on $\mathfrak{A}$ with respect to $\tau$ if the definitions of the corresponding relations of $\mathfrak{A}/E$ do not depend on the choices of the representatives of the equivalence classes of $E$: for every $R\in\tau$ and all tuples $\bar{t}$ we have $q_{E}(\bar{t})\in R^{\mathfrak{A}/{E}}$ if and only if $\bar{t}\in R^{\mathfrak{A}}$.
Next, recall from the introduction that an $\omega$-categorical structure $\mathfrak{A}$ is a model-complete core if the endomorphisms of $\mathfrak{A}$ preserve all orbits of $\mathrm{Aut}(\mathfrak{A})$. A reformulation of this definition that will be more convenient for us is that for every $e\in\mathrm{End}(\mathfrak{A})$ and every finite $F\subseteq A$, there exists $\alpha\in\mathrm{Aut}(\mathfrak{A})$ such that $e|_{F}=\alpha|_{F}$ [15]; we will use this in the proof of Theorem 2.2.
Now, we properly define the Ramsey property. For structures $\mathfrak{A}$ and $\mathfrak{W}$, we denote by $\binom{\mathfrak{W}}{\mathfrak{A}}$ the set of all embeddings of $\mathfrak{A}$ into $\mathfrak{W}$. A class $\mathcal{K}$ of structures over a common signature $\tau$ has the Ramsey property (RP) if, for all $\mathfrak{X},\mathfrak{Y}\in\mathcal{K}$ and $k\in\mathbb{N}$, there exists $\mathfrak{W}\in\mathcal{K}$ such that, for every map $f\colon\binom{\mathfrak{W}}{\mathfrak{X}}\rightarrow[k]$, there exists $e\in\binom{\mathfrak{W}}{\mathfrak{Y}}$ such that $f$ is constant on the set $\{e\circ u\;|\;u\in\textstyle\binom{\mathfrak{Y}}{\mathfrak{X}}\}\subseteq\textstyle\binom{\mathfrak{W}}{\mathfrak{X}}.$ It is a folklore fact that considering the special case $k=2$ is enough for verifying the RP [49], and we call any $\mathfrak{W}$ satisfying the above in this case a Ramsey witness for $(\mathfrak{X},\mathfrak{Y})$. A homogeneous structure is Ramsey if its age has the Ramsey property.
A.2. The KPT correspondence.
In the proof of Proposition 4.5, we use the Kechris-Pestov-Todorčević (KPT) correspondence, which links the Ramsey property for homogeneous structures over countable signatures to a property of their automorphism group viewed as a topological group.
A topological group is a pair $(G,\mathcal{T})$, where $G$ is a group and $(G;\mathcal{T})$ a topological space, such that both taking inverses and composition in the group $G$ are continuous with respect to $\mathcal{T}$. A topological group $(G,\mathcal{T})$ is called extremely amenable if every continuous action $\alpha$ of $(G,\mathcal{T})$ on a topological compact Hausdorff space $(X,\mathcal{T}^{\prime})$ has a fixed point, i.e. there exists $x\in X$ such that $\alpha(g)(x)=x$ for every $g\in G$; the action $\alpha$ being continuous means that the map $(g,x)\mapsto\alpha(g)(x)$ is. A bijective map $f$ between two topological groups is a topological isomorphism if it is a group isomorphism and both $f$ and $f^{-1}$ are continuous. Clearly, topological group isomorphisms preserve extreme amenability.
The formulation of the KPT correspondence below (Theorem A.1) refers to the extreme amenability of $\mathrm{Aut}(\mathfrak{B})$, without mentioning any specific topology: this is because it is understood to be equipped with the topology of pointwise convergence. This is the smallest topology on $\mathrm{Aut}(\mathfrak{B})$ containing all sets of the form $\{g\in\mathrm{Aut}(\mathfrak{B})\mid g(\bar{t})=\bar{s}\}$ for some tuples $\bar{t},\bar{s}$ over $B$, the basic open sets of the topology.
Theorem A.1 (Kechris, Pestov, and Todorcevic [52]).
Let $\mathfrak{B}$ be a homogeneous structure over a countable relational signature. Then the following are equivalent:
  • $\mathrm{Aut}(\mathfrak{B})$ is extremely amenable;
  • $\mathrm{age}(\mathfrak{B})$ has the RP.
A.3. A proof of Proposition 4.1
See 4.1
Proof.
Let $\mathfrak{B}$ be a CSP-injective $\omega$-categorical model-complete core. We show that $\mathfrak{B}$ has no algebraicity. Suppose, on the contrary, that there exists $\bar{b}\in B^{k}$ and $b^{\prime}\notin\bar{b}$ such that the unary relation $\{b^{\prime}\}$ is first-order definable in $\mathfrak{B}$ using $\bar{b}$ as parameters. By [15, Theorem 4.5.1 and Lemma 4.5.5], the expansion $(\mathfrak{B},\{\bar{b}\})$ is an $\omega$-categorical model-complete core, and there exists a primitive positive formula $\phi(x)$ defining $\{b^{\prime}\}$ in $(\mathfrak{B},\{\bar{b}\})$. Also, by [15, Theorem 4.5.1], there exists a primitive positive formula $\psi(\bar{y})$ defining the orbit of $\bar{b}$ in $\mathfrak{B}$. We may assume that $\psi$ does not contain any equality atoms, otherwise we eliminate them by identifying variables. Let $\phi^{\prime}$ be the primitive positive formula obtained by replacing the constants $\bar{b}$ in $\phi$ with the variables $\bar{y}$, and set $\phi^{\prime\prime}\coloneqq\psi(\bar{y})\wedge\phi^{\prime}(x_{1})\wedge\phi^{\prime}(x_{2})$; we may again assume that $\phi^{\prime}$ and $\phi^{\prime\prime}$ do not contain any equality atoms. We can use $\phi^{\prime\prime}$ to create an instance $\mathfrak{A}$ of $\mathrm{CSP}(\mathfrak{B})$ that maps homomorphically to $\mathfrak{B}$ but has no injective such homomorphism. More specifically, the domain of $\mathfrak{A}$ consists of the variables of $\phi^{\prime\prime}$, and the relations are determined by the atomic subformulas of $\phi^{\prime\prime}$. Clearly, $\mathfrak{A}$ has a homomorphism to $\mathfrak{B}$, because we can substitute $\bar{b}$ for $\bar{y}$ in $\phi^{\prime\prime}$ and $b$ for both $x_{1}$ and $x_{2}$. However, $\mathfrak{A}$ has no injective homomorphism to $\mathfrak{B}$. Indeed, if this was the case, then there would exist distinct witnesses $\bar{y}$, $x_{1}$, and $x_{2}$ for the satisfaction of $\phi^{\prime\prime}$ in $\mathfrak{B}$. By definition, there would exist $\alpha\in\mathrm{Aut}(\mathfrak{B})$ with $\alpha(\bar{y})=\bar{b}$. But then $\alpha(x_{1})=\alpha(x_{2})=c$, a contradiction to the injectivity of $\alpha$. We conclude that $\mathfrak{A}$ does not have any injective homomorphism to $\mathfrak{B}$, and hence $\mathfrak{B}$ is not CSP-injective, a contradiction to our original assumption. Hence, $\mathfrak{B}$ has no algebraicity. ∎
A.4. A proof of Proposition 4.2
See 4.2
Proof.
We instead show that a $\omega$-categorical $\mathrm{Pol}$-injective structure cannot be CSP-injective. Note that this leads to the same conclusion as the original statement, namely, that the formal intersection between the two properties is empty.
Let $\mathfrak{B}$ be an $\omega$-categorical Pol-injective structure. Note that every essentially injective operation preserves the relation $I_{4}$. Hence, every polymorphism of $\mathfrak{B}$ preserves $I_{4}$, and so $I_{4}$ is pp-definable in $\mathfrak{B}$. Let $\phi(x,y,u,v)$ be a primitive positive formula that defines $I_{4}$ in $\mathfrak{B}$. Since $I_{4}$ contains all $4$-tuples with pairwise distinct entries, we may assume that $\phi$ contains no equality atoms, otherwise we eliminate them by identifying variables. But now we can use $\phi$ to create an instance $\mathfrak{A}$ of $\mathrm{CSP}(\mathfrak{B})$ that maps homomorphically to $\mathfrak{B}$ but has no injective such homomorphism. More specifically, the domain of $\mathfrak{A}$ consists of the variables of $\phi$, except that we identify $x$ and $y$ (but not $u$ and $v$), and the relations are determined by the atomic subformulas of $\phi$. We conclude that $\mathfrak{B}$ is not CSP-injective. ∎
A.5. A proof of Proposition 4.3
See 4.3
Proof.
The Pol-injectivity of $\mathfrak{A}_{I_{4}}$ follows from the known fact that every polymorphism of $I_{4}$ is essentially injective [16] (see also [15, Lemma 7.5.1]). We proceed with the proof of Datalog-interreducibility.
We first give a Datalog-reduction $\mathcal{I}^{\ast}$ from $\mathrm{CSP}(\mathfrak{A}_{I_{4}})$ to $\mathrm{CSP}(\mathfrak{A})$. Let $\tau$ be the signature of $\mathfrak{A}$. Given an instance $\mathfrak{J}$ of $\mathrm{CSP}(\mathfrak{A}_{I_{4}})$, the relation $\approx$ is defined in $\mathfrak{J}$ by the following Datalog program:
$(x\approx y)\Leftarrow$ $I_{4}(u,v,x,y)\wedge(u\approx v),$
$(x\approx z)\Leftarrow$ $(x\approx y)\wedge(y\approx z),$
$(x\approx y)\Leftarrow$ $(y\approx x),$
$(x\approx y)\Leftarrow$ $(x=y).$
[…]
Suppose that $\mathcal{I}^{\ast}(\mathfrak{J})\rightarrow\mathfrak{A}$. Note that $(\mathfrak{J}/{\approx})|^{\tau}$ is a substructure of $\mathcal{I}^{\ast}(\mathfrak{J})$, and hence $(\mathfrak{J}/{\approx})|^{\tau}\rightarrow\mathfrak{A}$. Since $\mathfrak{A}$ is CSP-injective, there exists an injective homomorphism $h\colon(\mathfrak{J}/{\approx})|^{\tau}\rightarrow\mathfrak{A}$. By the definition of $\approx$, if $([s_{1}]_{\approx},[s_{2}]_{\approx},[s_{3}]_{\approx},[s_{4}]_{\approx})\in I_{4}^{\mathfrak{J}/{\approx}}$ and $[s_{1}]_{\approx}=[s_{2}]_{\approx}$, then $[s_{3}]_{\approx}=[s_{4}]_{\approx}$. Since $h$ is injective, it follows that $h$ is also a homomorphism from $\mathfrak{J}/{\approx}$ to $\mathfrak{A}_{I_{4}}$. Then $h\circ q_{\approx}$ is a homomorphism from $\mathfrak{J}$ to $\mathfrak{A}_{I_{4}}$.
Now suppose that there exists a homomorphism $g\colon\mathfrak{J}\rightarrow\mathfrak{A}_{I_{4}}$. Note that, by definition, for every $k$-ary $R\in\tau$, we have $(s_{i_{1}},\dots,s_{i_{k}})\in R^{\mathcal{I}^{\ast}(\mathfrak{J})}$ if and only if there are $j_{1}\in[s_{i_{1}}]_{\approx},\dots,j_{k}\in[s_{i_{k}}]_{\approx}$ such that $(j_{1},\dots,j_{k})\in R^{\mathfrak{J}}$. By the definition of $\approx$, every function preserving $I_{4}$ must identify $\approx$-equivalent elements, and therefore $(g(s_{i_{1}}),\dots,g(s_{i_{k}}))=(g(j_{1}),\dots,g(j_{k}))$, which implies $(g(s_{i_{1}}),\dots,g(s_{i_{k}}))\in R^{\mathfrak{A}_{I_{4}}}$. Therefore, $g$ is a homomorphism from $\mathcal{I}^{\ast}(\mathfrak{J})$ to $\mathfrak{A}$.
The reduction in the opposite direction is trivial. Given an instance $\mathfrak{J}$ of $\mathrm{CSP}(\mathfrak{A})$, we define $\mathcal{I}(\mathfrak{J})$ as the $(\tau\cup\{I_{4}\})$-expansion of $\mathfrak{J}$ by an empty relation. ∎
A.6. A proof of Proposition 4.5
See 4.5
Proof.
For (1), we first show that $\mathrm{Aut}(\mathfrak{A})\wr\mathrm{Aut}(\mathfrak{B})$ is a subgroup of $\mathrm{Aut}(\mathfrak{A}\wr\mathfrak{B})$. Clearly, the wreath product is a subgroup of $\mathrm{Sym}(A\times B)$ and preserves $E^{\mathfrak{A}\wr\mathfrak{B}}$. Moreover, as every group element $\left((g_{b})_{b\in B},h\right)$ simply acts like $h$ in the second component, all $\sigma$-relations are preserved. Concerning $\tau$-relations, note that all pairs occurring within a tuple $((a_{1},b_{1}),\dots,(a_{k},b_{k}))\in R^{\mathfrak{A}\wr\mathfrak{B}}$ share the same second entry $b=b_{1}=\cdots=b_{k}$. Therefore, the same $g_{h(b)}\in\mathrm{Aut}(\mathfrak{A}))$ is picked to act on all the first entries, and the image $((g_{h(b)}(a_{1}),h(b)),\dots,(g_{h(b)}(a_{k}),h(b)))$ lies within $R^{\mathfrak{A}\wr\mathfrak{B}}$.
To prove the equality of the two groups, it remains to show that every $w\in\mathrm{Aut}(\mathfrak{A}\wr\mathfrak{B})$ can be written in the above form. As $w$ acts on the equivalence classes of $E^{\mathfrak{A}\wr\mathfrak{B}}$, it induces an automorphism $h^{\prime}$ on the $\sigma$-reduct of $(\mathfrak{A}\wr\mathfrak{B})/E^{\mathfrak{A}\wr\mathfrak{B}}$, which is isomorphic to $\mathfrak{B}$. Further, for any $b\in B$, $w$ also induces an automorphism $g^{\prime}$ on the $\tau$-reduct of the substructure induced by $A\times\{b\}$, which is in turn isomorphic to $\mathfrak{A}$. Denoting the corresponding automorphisms of $\mathfrak{A}$ and $\mathfrak{B}$ by $g_{h(b)},b\in B$ and $h$, it is now easy to check that $((g_{b})_{b\in B},h)$ and $w$ coincide.
For (2), we must show that fixing any tuple $\bar{t}$ of parameters does not yield any nontrivial fixed point of the automorphism group when stabilizing $\bar{t}$. Let $(a^{\prime},b^{\prime})$ be any element outside $\bar{t}$. Let $A^{\prime}$ be the set of all elements $a\in A$ such that $(a,b)$ appears in $\bar{t}$ for some $b\in B$. Move $a^{\prime}$ to a different element via some $g\in\mathrm{Aut}(\mathfrak{A})$ whilst fixing $A^{\prime}$; this is possible because $\mathfrak{A}$ has no algebraicity. Then $\left((g_{b})_{b\in B},id\right)$ with $g_{b}\coloneqq g$ if $b=b^{\prime}$ and $id$ otherwise is an automorphism of $\mathrm{Aut}(\mathfrak{A}\wr\mathfrak{B})$ fixing every entry in $\bar{t}$ and moving $(a^{\prime},b^{\prime})$.
For (3), recall that, by item (1), we have $\mathrm{Aut}(\mathfrak{A}\wr\mathfrak{B})=\mathrm{Aut}(\mathfrak{A})\wr\mathrm{Aut}(\mathfrak{B})$. But note that $\mathrm{Aut}(\mathfrak{A})\wr\mathrm{Aut}(\mathfrak{B})$ in particular contains the Cartesian product $\mathrm{Aut}(\mathfrak{A})\times\mathrm{Aut}(\mathfrak{B})$. Since the Cartesian product already witnesses that there are only finitely many orbits of tuples of each arity, the statement follows.
For (4), let $\bar{t}$ and $\bar{s}$ be tuples of the same length such that the map sending $\bar{t}$ to $\bar{s}$ is a partial isomorphism. By reordering, we split $\bar{t}$ into subtuples $(\bar{t}_{b_{1}},\dots,\bar{t}_{b_{n}})$ for distinct elements $b_{i}\in B$ such that the second projection of all entries of $\bar{t}_{b_{i}}$ is $b_{i}$. Similarly, we obtain $(\bar{s}_{c_{1}},\dots,\bar{s}_{c_{n}})$. Pick $g_{c_{i}}\in\mathrm{Aut}(\mathfrak{A})$ sending the first projections of $\bar{t}_{b_{i}}$ to those of $\bar{s}_{c_{i}}$. Pick $h\in\mathrm{Aut}(\mathfrak{B})$ sending $b_{i}$ to $c_{i}$. The automorphism $((g_{b})_{b\in B}),h)$, where $g_{b}=id$ if hitherto undefined, sends $\bar{t}$ to $\bar{s}$.
For (5), let $\Phi_{\mathfrak{A}}$ and $\Phi_{\mathfrak{B}}$ be any universal sentences defining $\mathrm{age}(\mathfrak{A})$ and $\mathrm{age}(\mathfrak{B})$, respectively. Denote their quantifier-free parts by $\phi_{\mathfrak{A}}$ and $\phi_{\mathfrak{B}}$. For two tuples $\bar{x}_{1},\bar{x}_{2}$ of the same arity $k$, we use $E(\bar{x}_{1},\bar{x}_{2})$ as a shortcut for the formula stating that, for every $i\in[k]$, the $i$-the entry of $\bar{x}_{1}$ and the $i$-the entry of $\bar{x}_{2}$ are contained in an $E$-atom. Then the following sentence $\Phi_{\mathfrak{A}}\wr\Phi_{\mathfrak{B}}$ defines $\mathrm{age}(\mathfrak{A}\wr\mathfrak{B})$:
$\forall x,y,z\Big{(}E(x,x)\wedge\big{(}E(x,y)\Rightarrow E(y,x)\big{)}\wedge\big{(}E(x,y)\wedge E(y,z)\Rightarrow E(x,z)\big{)}$
${}\wedge\bigwedge\nolimits_{R\in\sigma}\forall\bar{x}_{1},\bar{x}_{2}\Big{(}E(\bar{x}_{1},\bar{x}_{2})\Rightarrow\big{(}R(\bar{x}_{1})\Leftrightarrow R(\bar{x}_{2})\big{)}\Big{)}$
${}\wedge\forall\bar{x}\Big{(}\phi_{\mathfrak{B}}(\bar{x})\vee\bigvee\nolimits_{x_{1},x_{2}\in\bar{x}}\big{(}E(x_{1},x_{2})\wedge(x_{1}\neq x_{2})\big{)}\Big{)}$
${}\wedge\bigwedge\nolimits_{R\in\tau}\forall\bar{x}\Big{(}R(\bar{x})\Rightarrow\bigwedge\nolimits_{x_{1},x_{2}\in\bar{x}}E(x_{1},x_{2})\Big{)}$
${}\wedge\forall\bar{x}\Big{(}\phi_{\mathfrak{A}}(\bar{x})\vee\bigvee\nolimits_{x_{1},x_{2}\in\bar{x}}\neg E(x_{1},x_{2})\Big{)}\Big{)}.$
To see this, note that:
  • the first line ensures that $E$ is an equivalence relation,
  • the second line ensures that $E$ is a relational congruence w.r.t. $\sigma$-relations,
  • the third line ensures that the factor by $E$ satisfies $\Phi_{\mathfrak{B}}$,
  • the fourth line ensures that $\tau$-relations do not stretch over different $E$-classes,
  • the fifth line ensures that each $E$-class satisfies $\Phi_{\mathfrak{A}}$.
For (6), recall that, by Theorem A.1, $\mathrm{Aut}(\mathfrak{A})$ and $\mathrm{Aut}(\mathfrak{B})$ are extremely amenable. By Proposition 4.5(1), we have $\mathrm{Aut}(\mathfrak{A}\wr\mathfrak{B})=\mathrm{Aut}(\mathfrak{A})\wr\mathrm{Aut}(\mathfrak{B})$. Hence, we can use the folklore fact that $\mathrm{Aut}(\mathfrak{A})\wr\mathrm{Aut}(\mathfrak{B})$ is a semidirect product $\mathrm{Aut}(\mathfrak{A})^{B}\rtimes\mathrm{Aut}(\mathfrak{B})$. In particular, we immediately get that $\mathrm{Aut}(\mathfrak{A})^{B}$ is a normal subgroup of $\mathrm{Aut}(\mathfrak{A}\wr\mathfrak{B})$. We claim that $\mathrm{Aut}(\mathfrak{A})^{B}$ is closed with respect to the topology of pointwise convergence. This follows directly from the fact that the complement of $\mathrm{Aut}(\mathfrak{A})^{B}$ can be written as
$\bigcup\nolimits_{x\in B,y\in B\setminus\{x\}}\{\left((g_{b})_{b\in B},h)\right)\in\mathrm{Aut}(\mathfrak{A}\wr\mathfrak{B})\mid h(x)=y\},$
[…]
A.7. A full proof of Theorem 2.2 and Theorem 2.4
See 2.2
See 2.4
As the two theorems only differ in the presence of the relation $I_{4}$, large parts of the proofs of Theorem 2.2 and Theorem 2.4 are essentially the same. For this reason, we conduct both proofs simultaneously, and differentiate whenever necessary.
Proof of Theorem 2.2 and Theorem 2.4.
Let $\tau$ and $\sigma$ be the signatures of $\mathfrak{A}$ and $\mathfrak{B}$, respectively. We first define the auxiliary structure $\mathfrak{B}^{\prime}\coloneqq(\mathbb{Q};<)\wr\mathfrak{B}$. By Example 1.3 and Example 2.1, the structure $(\mathbb{Q};<)$ is finitely bounded, homogeneous, and has no algebraicity. It follows immediately from Proposition 4.5(2) that $\mathfrak{B}^{\prime}$ has no algebraicity. The expansion $\mathfrak{B}^{\prime}_{\neq}$ of $\mathfrak{B}^{\prime}$ by the binary inequality has no algebraicity because taking expansions by first-order definable relations does not change the automorphism group; the same is true for the structure $\mathfrak{B}^{\prime}_{\neq,I_{4}}$. Note that removing all symbols in $\sigma\setminus\tau$ as well as the symbol $<$ from $\mathfrak{B}^{\prime}_{\neq}=\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq}$ yields the structure $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. We continue with the proofs of CSP/Pol-injectivity and Datalog-interreducibility, starting with the CSP-injective case.
We claim that $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ is CSP-injective. To this end, let $\mathfrak{J}$ be an arbitrary finite structure for which there exists a homomorphism $h\colon\mathfrak{J}\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. We define $\mathfrak{J}_{\pi}$ as the $(\tau\cup\{E\})$-structure with domain $J$ such that, for every $R\in\tau\cup\{E\}$, we have $\bar{t}\in R^{\mathfrak{J}_{\pi}}$ if and only if $h(\bar{t})\in R^{(\mathbb{Q}\wr\mathfrak{A})_{\neq}}$. It is easy to see that $\mathfrak{J}_{\pi}$ embeds into the $(\tau\cup\{E\})$-reduct of $\mathfrak{B}^{\prime}$ (note that if multiple elements in $J$ are mapped to $(q,a)$ for some $q\in\mathbb{Q}$ and $a\in A$, we can always map them to different elements in ×$\mathbb{Q}\times\{a\}$ since this does not affect the relations in $\tau\cup\{E\}$). Hence, $\mathfrak{J}_{\pi}$ expanded by the binary inequality predicate embeds into $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ and admits an injective homomorphism from $\mathfrak{J}$. We conclude that $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ is CSP-injective.
Next, we informally describe two Datalog-reductions $\mathcal{I}^{\ast}$ and $\mathcal{I}$: from $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$ to $\mathrm{CSP}(\mathfrak{A})$ and back, starting with the former. Let $\mathfrak{J}$ be an instance of $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$. Denote by $\sim$ the equivalence closure of $E^{\mathfrak{J}}$, and let $q_{\sim}\colon\mathfrak{J}\rightarrow(\mathfrak{J}/{\sim}),x\mapsto[x]_{\sim}$ be the induced quotient map. We define $\mathcal{I}^{\ast}(\mathfrak{J})$ as the $\tau$-structure with domain $J$ and whose relations are the preimages of all $\tau$-relations in $\mathfrak{J}/{\sim}$ under $q_{\sim}$, except that we additionally add, for every $x\in J$ such that $(x,x)\in{\neq}^{\mathfrak{J}}$, the tuple $(x,\dots,x)$ to every $\tau$-relation. It is easy to see that $\mathcal{I}^{\ast}(\mathfrak{J})$ is definable in Datalog; we verify that $\mathcal{I}^{\ast}$ is a reduction from $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$ to $\mathrm{CSP}(\mathfrak{A})$. Clearly, if $\mathfrak{J}$ maps homomorphically to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$, then $\mathcal{I}^{\ast}(\mathfrak{J})$ maps homomorphically to $\mathfrak{A}$. Suppose that there exists a homomorphism $h\colon\mathcal{I}^{\ast}(\mathfrak{J})\rightarrow\mathfrak{A}$. Since by non-triviality $\mathfrak{A}$ does not have a constant endomorphism, there is no $x\in J$ such that $(x,x)\in{\neq}^{\mathfrak{J}}$: otherwise, the map on $\mathfrak{A}$ sending every element to $h(q_{\sim}(x))$ would provide such constant endomorphism. Denote by $\mathfrak{J}^{\prime}$ the structure obtained from $\mathfrak{J}$ by removing all tuples from ${\neq}^{\mathfrak{J}}$. Now, choosing representatives $j_{1},\dots,j_{k}$ for all classes of $\mathfrak{J}/{\sim}$, the function sending the entirety of $[j_{i}]_{\sim}$ to $(0,h(j_{i}))$ for all $i\leq k$ is a homomorphism from $\mathfrak{J}^{\prime}$ to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. By the CSP-injectivity of $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$, it can be replaced by an injective homomorphism, which is then also a homomorphism from $\mathfrak{J}$ to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. We continue with the reduction from $\mathrm{CSP}(\mathfrak{A})$ to $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$, which is trivial. Let $\mathfrak{J}$ be an instance of $\mathrm{CSP}(\mathfrak{A})$. We define $\mathcal{I}(\mathfrak{J})$ as the $(\tau\cup\{E,\neq\})$-expansion of $\mathfrak{J}$ by empty relations. By CSP-injectivity, there exists a homomorphism $h\colon\mathcal{I}(\mathfrak{J})\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ if and only if there exists an injective such homomorphism $i$. This is the case if and only if the relation $E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq}}$ restricted to $i(J)$ describes the kernel of a homomorphism from $\mathfrak{J}$ to $\mathfrak{A}$. Therefore, $\mathfrak{J}\rightarrow\mathfrak{A}$ if and only if $\mathcal{I}(\mathfrak{J})\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$.
We now consider $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$. By Proposition 4.3, $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ is Pol-injective and $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}})$ and $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$ are Datalog-interreducible. That $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}})$ and $\mathrm{CSP}(\mathfrak{A}$) are Datalog-interreducible then follows from the fact that Datalog-reductions can be composed.
Finally, we verify the properties in items 123, and 4. We only explicitly cover the case where $I_{4}$ is present, it is easy to see that our arguments also hold in the case without $I_{4}$.
Proof of item 1. The first part follows directly from Proposition 4.5(3) applied to $\mathfrak{B}^{\prime}$ and the fact that we take an expansion by relations which are preserved by all bijections, and in particular all automorphisms of $\mathfrak{B}^{\prime}$. We continue with the second part.
Suppose that $e^{\ast}$ is an endomorphism of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$. Let $S^{\ast}$ be an arbitrary finite subset of ×$\mathbb{Q}\times A$. Since $e^{\ast}$ preserves $\neq$, it is injective. Since $e^{\ast}$ preserves $E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}$, which is an equivalence relation, it naturally acts on its classes. Let $e$ be the map on (×$(\mathbb{Q}\times A)/E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}$ mapping each $[x]_{E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}}$ to $[e^{\ast}(x)]_{E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}}$, and let $S\coloneqq\{[x]_{E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}}\mid x\in S^{\ast}\}$. Similarly as in Proposition 4.5(1), we have that $e$ is an endomorphism of $\mathfrak{A}$. Since $\mathfrak{A}$ is a model-complete core, there exists an automorphism $\alpha$ of $\mathfrak{A}$ such that $\alpha|_{S}=e|_{S}$. We claim that there exists an automorphism ${\alpha}^{\ast}$ of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ such that ${\alpha}^{\ast}|_{S^{\ast}}=e^{\ast}|_{S^{\ast}}$. To see this, note that we can simply fix a permutation of each equivalence class of $E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}$ and extend $\alpha$, which is a permutation of the equivalence classes, according to it. Since $S^{\ast}$ is finite, we can moreover select the permutations so that ${\alpha}^{\ast}$ coincides with $e^{\ast}$ on $S^{\ast}$.
Next, suppose that $e$ is an endomorphism of $\mathfrak{A}$, and let $S$ be an arbitrary finite subset of $A$. Define $S^{\ast}\coloneqq\{0\}\times S$. Let $e^{\ast}$ be the map on ×$\mathbb{Q}\times A$ defined via $e^{\ast}(q,a)\coloneqq(q,e(a))$. This function is an endomorphism of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$. Since $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ is a model-complete core, there exists an automorphism ${\alpha}^{\ast}$ of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ such that ${\alpha}^{\ast}|_{S^{\ast}}=e^{\ast}|_{S^{\ast}}$. Let $\alpha$ be the map on $A$ mapping each $[x]_{E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}}$ to $[e^{\ast}(x)]_{E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}}$ As before, we have that $\alpha$ is a well-defined automorphism of $\mathfrak{A}$ extending $e$ on $S$.
Proof of item 2. Since $\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq,I_{4}}$ is an expansion of $\mathfrak{B}^{\prime}$ by two relations that are preserved by all embeddings, we can ignore these relations and only prove the statement for $\mathfrak{B}^{\prime}$. Recall that $\mathfrak{B}^{\prime}=(\mathbb{Q};<)\wr\mathfrak{B}$. The first part of item 2 follows directly from Proposition 4.5(4) because $(\mathbb{Q};<)$ is homogeneous. We continue with the second part. There the backward direction follows directly from Proposition 4.5(6) because $(\mathbb{Q};<)$ is homogeneous Ramsey [52]. We continue with the forward direction, which we prove by hand.
Suppose that $\mathrm{age}(\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq,I_{4}})$ has the RP. Then clearly also $\mathrm{age}(\mathfrak{B}^{\prime})$ has the RP. Let $\mathfrak{X},\mathfrak{Y}\in\mathrm{age}(\mathfrak{B})$ be arbitrary. We transform them into $\mathfrak{X}^{\prime},\mathfrak{Y}^{\prime}\in\mathrm{age}(\mathfrak{B}^{\prime})$ by letting $E$ interpret as the diagonal relation and $<$ as the empty relation. Since $\mathrm{age}(\mathfrak{B}^{\prime})$ has the RP, there exists a Ramsey witness $\mathfrak{W}^{\prime}\in\mathrm{age}(\mathfrak{B}^{\prime})$ for $(\mathfrak{X}^{\prime},\mathfrak{Y}^{\prime})$. We set $\mathfrak{W}\coloneqq(\mathfrak{W}^{\prime}/E^{\mathfrak{W}^{\prime}})|^{\sigma}$. Let $f\colon\binom{\mathfrak{W}}{\mathfrak{X}}\rightarrow[2]$ be arbitrary. We define $f^{\prime}\colon\binom{\mathfrak{W}^{\prime}}{\mathfrak{X}^{\prime}}\rightarrow[2]$ as follows. For an embedding $e^{\prime}\in\binom{\mathfrak{W}^{\prime}}{\mathfrak{X}^{\prime}}$, we set $f^{\prime}(e^{\prime})\coloneqq f(q_{E}\circ e^{\prime})$, where $q_{E}$ denotes the factor map $x\mapsto[x]_{E}$. We have that $q_{E}\circ e^{\prime}\in\binom{\mathfrak{W}}{\mathfrak{X}}$ because $E$ interprets as the diagonal relation in $\mathfrak{X}^{\prime}$ and as a relational congruence w.r.t. $\sigma$ in $\mathfrak{W}^{\prime}$. Since $\mathfrak{W}^{\prime}$ is a Ramsey witness for $(\mathfrak{X}^{\prime},\mathfrak{Y}^{\prime})$, there exists $e^{\prime}\in\binom{\mathfrak{W}^{\prime}}{\mathfrak{Y}^{\prime}}$ such that $f^{\prime}$ is constant on $\bigl{\{}e^{\prime}\circ u^{\prime}\;|\;u^{\prime}\in\binom{\mathfrak{Y}^{\prime}}{\mathfrak{X}^{\prime}}\bigr{\}}\subseteq\binom{\mathfrak{W}^{\prime}}{\mathfrak{X}^{\prime}}$. Since $E$ interprets as the diagonal relation in $\mathfrak{Y}^{\prime}$ and as a relational congruence w.r.t. $\sigma$ in $\mathfrak{W}^{\prime}$, the map $e\coloneqq q_{E}\circ e^{\prime}$ is contained in $\binom{\mathfrak{W}}{\mathfrak{Y}}$. By construction, $f$ is constant on $\bigl{\{}e\circ u\;|\;u\in\binom{\mathfrak{Y}}{\mathfrak{X}}\bigr{\}}\subseteq\binom{\mathfrak{W}}{\mathfrak{X}}$.
Proof of item 3. The finite boundedness can be deduced from Proposition 4.5(5), and the fact that we take an expansion by relations definable by a Boolean combination of equality atoms. More specifically, as explained in the introduction, up to isomorphism, the finite substructures of a finitely bounded structures are just the finite models of a universal sentence. By adding new clauses, that define the relations $\neq$ and $I_{4}$ in terms of equalities, to the universal sentence for $\mathrm{age}(\mathfrak{B}^{\prime})$, we obtain a universal sentence for $\mathrm{age}(\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq})$.
Proof of item 4. One direction is simple (and does not use the assumption on the orbit growth), because $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ clearly pp-constructs $\mathfrak{A}$; the pp-construction is given by $\mathfrak{A}=((\mathbb{Q}\wr\mathfrak{A})_{\neq}/E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq}})|^{\tau}$. Thus, if $\mathfrak{A}$ pp-constructs $\mathfrak{K}_{3}$, then so does $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$.
For the other direction, we use a characterization via the pseudo-Siggers identity from [7, Theorems 1.3 and 3.4]. Suppose that $\mathfrak{A}$ does not pp-construct $\mathfrak{K}_{3}$. Let $\mathfrak{C}$ be the model-complete core of $\mathfrak{A}$. Since $\mathfrak{A}$ and $\mathfrak{C}$ are homomorphically equivalent, $\mathfrak{C}$ also does not pp-construct $\mathfrak{K}_{3}$. Then, by Theorem 1.3 in [7], there exist $\alpha,\beta,s\in\mathrm{Pol}(\mathfrak{C})$ witnessing the pseudo-Siggers identity. Consider the structure $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$. By item 1, $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$ is an $\omega$-categorical model-complete core. Note that we can lift the homomorphic equivalence between $\mathfrak{A}$ and $\mathfrak{C}$ to a homomorphic equivalence between $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ and $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$ by setting, for any homomorphism $h$ between $\mathfrak{A}$ and $\mathfrak{C}$ and irrespectively of its direction, $h^{\ast}(x,y)\coloneqq(e(x,y),h(y))$, where $e$ is an arbitrary injection from the Cartesian product of $\mathbb{Q}$ with the domain of $h$ into $\mathbb{Q}$. The inclusion of the injection in the first coordinate ensures that $h^{\ast}$ preserves $I_{4}$ and $\neq$ (because they are preserved by all injective operations). Since $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ and $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$ are homomorphically equivalent, $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$ is the model-complete core of $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. Next, we lift the pseudo-Siggers identity from $\mathrm{Pol}(\mathfrak{C})$ to $\mathrm{Pol}((\mathbb{Q}\wr\mathfrak{C})_{\neq})$. Let $e$ be an arbitrary injection from ×$\mathbb{Q}\times C$ into $\mathbb{Q}$. The structure $(\mathbb{Q};\neq,I_{4})$ is an $\omega$-categorical model-complete core that does not pp-construct $\mathfrak{K}_{3}$ [15, Theorems 12.0.1, 12.7.3, 12.9.2, and Corollary 6.4.4]. Hence, by Theorem 1.3 in [7] (since pp-constructions are more general than pp-interpretations with parameters), there are $\alpha^{\prime},\beta^{\prime},s^{\prime}\in\mathrm{Pol}(\mathbb{Q};\neq,I_{4})$ witnessing the pseudo-Siggers identity. We set $\alpha^{\ast}(x,y)\coloneqq(\alpha^{\prime}(x),\alpha(y))$, $\beta^{\ast}(x,y)\coloneqq(\beta^{\prime}(x),\beta(y))$, and
$s^{\ast}((x_{1},y_{1}),\dots,(x_{6},y_{6}))\coloneqq(s^{\prime}(e(x_{1},y_{1}),\dots,e(x_{6},y_{6})),s(y_{1},\dots,y_{6})).$
It is easy to verify that $\alpha^{\ast},\beta^{\ast},s^{\ast}$ are polymorphisms of $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$ and that these operations witness the pseudo-Siggers identity. It moreover follows by inspection of the proof of Proposition 4.5(3) that also the structure $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ has less than double exponential orbit growth, and therefore so does its model-complete core $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$ by the proof of [7, Corollary 1.8]. Combining this with the pseudo-Siggers identity, it follows from Theorem 3.4 in [7] (whose item (ii) is equivalent to the pp-construction of $\mathfrak{K}_{3}$ by Theorem 1.8 in [9]) that $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ does not pp-construct $\mathfrak{K}_{3}$. ∎
A.8. A proof of Corollary 2.5
See 2.5
Proof.
Suppose that $\mathfrak{A}$ does not pp-construct $\mathfrak{K}_{3}$. By Theorem 2.4(4), the structure $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ does not pp-construct $\mathfrak{K}_{3}$ either. Let $\mathfrak{C}$ be the model-complete core of $\mathfrak{A}$. By the proof of Theorem 2.4(4), $(\mathbb{Q}\wr\mathfrak{C})_{\neq,I_{4}}$ is the model-complete core of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$. Hence, also $(\mathbb{Q}\wr\mathfrak{C})_{\neq,I_{4}}$ does not pp-construct $\mathfrak{K}_{3}$. By [55, Theorem 3.14], $\mathrm{Pol}((\mathbb{Q}\wr\mathfrak{C})_{\neq,I_{4}})$ contains a binary essential operation $f$. Let $g$ and $h$ be any homomorphisms witnessing the homomorphic equivalence between $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ and $(\mathbb{Q}\wr\mathfrak{C})_{\neq,I_{4}}$. Then $(x,y)\mapsto g\circ f(h(x),h(y))$ is a binary essential polymorphism of $(\mathbb{Q}\wr\mathfrak{C})_{\neq,I_{4}}$. Since $(\mathbb{Q}\wr\mathfrak{C})_{\neq,I_{4}}$ is Pol-injective, this operation must be injective. ∎
Appendix B $\omega$-categorical cheeses for PCSPs
B.1. A proof of Proposition 5.1
For an operation $f\colon A^{n}\rightarrow A$, we denote the component-wise action of $f$ on $d$-tuples by $f\curvearrowright A^{d}$. For a set $F$ of operations on $A$, we denote by $F\curvearrowright A^{d}$ the set $\{f\curvearrowright A^{d}\mid f\in F\}$. See 5.1
Proof.
For $i\in[d]$, we denote by $\iota_{i}$ the function from $[1]$ to $[d]$ defined by $\iota_{i}(1)\coloneqq i$.
For (1), let $f\in\mathrm{Pol}({\mathfrak{A}}^{[d]})$ be arbitrary, and let $n$ be its arity. In what follows, we shall use the abbreviation $\bar{a}=(a_{1},\dots,a_{n})$. If $\bar{a}$ appears as the $i$-th row of $(\bar{x}_{1},\dots,\bar{x}_{n})\in(A^{d})^{n}$ (viewed as a $d\times n$-matrix) and the $j$-th row in $(\bar{y}_{1},\dots,\bar{y}_{n})\in(A^{d})^{n}$, then, since $f$ preserves $S^{{\mathfrak{A}}^{[d]}}_{\iota_{i},\iota_{j}}$, the $i$-th entry of $f(\bar{x}_{1},\dots,\bar{x}_{n})$ equals the $j$-th entry of $f(\bar{y}_{1},\dots,\bar{y}_{n})$. This ensures that there exists a well-defined operation $f^{\prime}\colon A^{n}\to A$ such that $f=f^{\prime}\curvearrowright A^{d}$. The definition of $R^{{\mathfrak{A}}^{[d]}}_{\mathrm{id}}$ for every $k$-ary relation $R^{\mathfrak{A}}$, where $\mathrm{id}$ is the identity map on $[k]$, guarantees that $f^{\prime}\in\mathrm{Pol}(\mathfrak{A})$. Thus, $\mathrm{Pol}({\mathfrak{A}}^{[d]})\subseteq\mathrm{Pol}(\mathfrak{A})\curvearrowright A^{d}$. The other inclusion is trivial.
For (2), we proceed as in the proof of [15, Lemma 3.5.4]; the identity map yields a $d$-dimensional pp-interpretation of ${\mathfrak{A}}^{[d]}$ in $\mathfrak{A}$ and the first projection gives a 1-dimensional pp-interpretation of $\mathfrak{A}$ in ${\mathfrak{A}}^{[d]}.$
For (3), note that $\mathrm{Aut}({\mathfrak{A}}^{[d]})=\mathrm{Aut}(\mathfrak{A})\curvearrowright A^{d}$ by item (1). Since every $n$-orbit in ${\mathfrak{A}}^{[d]}$ corresponds to a $dn$-orbit in $\mathfrak{A}$, there are finitely many $n$-orbits in ${\mathfrak{A}}^{[d]}$ for each $n\in\mathbb{N}$. Hence, ${\mathfrak{A}}^{[d]}$ is $\omega$-categorical as well.
For (4), let $\alpha:\mathfrak{D}_{1}\to\mathfrak{D}_{2}$ be an isomorphism between two finite substructures of ${\mathfrak{A}}^{[d]}$ and let $A_{1},A_{2}$ be the sets of elements of $A$ appearing as entries in some $d_{1}\in D_{1}$ or $d_{2}\in D_{2}$, respectively. Then the relations $S^{{\mathfrak{A}}^{[d]}}_{\iota_{i},\iota_{j}}$ and $\widehat{R}^{\,{\mathfrak{A}}^{[d]}}$ for $R\in\tau$ guarantee that $\alpha$ induces an isomorphism $\beta$ between $\mathfrak{A}_{1}$ and $\mathfrak{A}_{2}$. Since $\mathfrak{A}$ is homogeneous, we can extend $\beta$ to some automorphism $\beta^{\ast}$ of $\mathfrak{A}$. Now $\alpha^{\ast}\coloneqq\beta^{\ast}\curvearrowright A^{d}$ is an automorphism of ${\mathfrak{A}}^{[d]}$ extending $\alpha$.
For (5), let $e\in\mathrm{End}({\mathfrak{A}}^{[d]})$ be arbitrary and let $D$ be an arbitrary finite subset of $A^{d}$. Define $A_{D}$ as the set of all elements of $A$ appearing in an entry of some $d\in D$. By item (1), we have $e=e^{\prime}\curvearrowright A^{d}$ for some $e^{\prime}\in\mathrm{End}(\mathfrak{A})$. Since $\mathfrak{A}$ is a model-complete core, there exists $\alpha^{\prime}\in\mathrm{Aut}(\mathfrak{A})$ such that $\alpha^{\prime}|_{A_{D}}=e^{\prime}|_{A_{D}}$. Then $\alpha\coloneqq\alpha^{\prime}\curvearrowright A^{d}$ is an automorphism of ${\mathfrak{A}}^{[d]}$ satisfying $\alpha|_{D}=e|_{D}$.
Item (6) can be shown in a similar fashion as [15, Lemma 3.5.4].
For (7), note that, by item (1) and since the topology of pointwise convergence is determined by the behaviour of functions on finite subsets, the mapping $\alpha\mapsto\alpha\curvearrowright A^{d}$ is a topological isomorphism from $\mathrm{Aut}(\mathfrak{A})$ to $\mathrm{Aut}({\mathfrak{A}}^{[d]})$. By Theorem A.1, we have that ${\mathfrak{A}}^{[d]}$ is homogeneous Ramsey.
Item (8) follows directly from the definition of ${\mathfrak{A}}^{[d]}$. ∎
B.2. A full proof of Proposition 5.2
See 5.2
We split the proof of Proposition 5.2 into several individual statements, which are either independent or build on top of each other (in the order in which they are presented). We start with the following lemma, a crucial independent observation about pseudo-cyclic polymorphisms from finite substructures of $\mathfrak{A}$ to $\mathfrak{A}$ itself.
Lemma B.1.
Let $\mathfrak{A}$ be a reduct of a linearly ordered finitely bounded homogeneous structure $\mathfrak{B}$ and let $\mathfrak{S}$ be a finite substructure of $\mathfrak{A}$. If $f\in\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ is pseudo-cyclic modulo $\mathrm{Aut}(\mathfrak{B})$, then $f$ is cyclic.
Proof.
Let $n$ be the arity of $f$ and let $\alpha_{1},\alpha_{2}\in\mathrm{Aut}(\mathfrak{B})$ be such that
$\alpha_{1}\circ f(s_{1},\dots,s_{n})=\alpha_{2}\circ f(s_{2},\dots,s_{n},s_{1})$
for all $s_{1},\dots,s_{n}\in S$. Suppose, on the contrary, that $f$ is not cyclic. Then there are $s_{1},\dots,s_{n}\in S$ such that $f(s_{1},\dots,s_{n})\neq f(s_{2},\dots,s_{n},s_{1})$. Let $<$ denote the linear order of $\mathfrak{B}$ and set $\tilde{s_{i}}\coloneqq(s_{i},\dots,s_{n},s_{1},\dots,s_{i-1})$. By assumption, we have $f(\tilde{s_{1}})\neq f(\tilde{s_{2}})$. Without loss of generality, $f(\tilde{s_{1}})<f(\tilde{s_{2}})$. Since $\alpha_{1},\alpha_{2}$ preserve $<$ and its complement, we have
$f(\tilde{s_{1}})<f(\tilde{s_{2}})\Leftrightarrow$ $\enspace\alpha_{1}(f(\tilde{s_{1}}))<\alpha_{1}(f(\tilde{s_{2}}))$
$\Leftrightarrow$ $\enspace\alpha_{2}(f(\tilde{s_{2}}))<\alpha_{2}(f(\tilde{s_{3}}))$
$\Leftrightarrow$ $\enspace f(\tilde{s_{2}})<f(\tilde{s_{3}})$
[…]
$\Leftrightarrow$ $\enspace f(\tilde{s_{n}})<f(\tilde{s_{1}}),$
a contradiction. Thus, $f$ is cyclic. ∎
The next lemma essentially states that $\mathcal{I}(\mathfrak{X})\coloneqq{\mathfrak{X}}^{[d]}$ is a reduction from $\mathrm{CSP}(\mathfrak{A})$ to $\mathrm{CSP}({\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$. We will, however, use it in the more general setting of PCSPs.
Lemma B.2.
Let $\mathfrak{A}$ be a reduct of a finitely bounded homogeneous structure $\mathfrak{B}$. Suppose that the bounds of $\mathfrak{B}$ have size $\leq d$ and its relations are of arity $\leq d-1$. Then, for every finite structure $\mathfrak{X}$ and every homomorphism $h\colon{\mathfrak{X}}^{[d]}\to{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$, there exists a homomorphism $g\colon\mathfrak{X}\to\mathfrak{A}$ such that $h(\bar{x})=g(\bar{x})_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$ for every $\bar{x}\in X^{d}$.
Proof.
Let $h$ and $\mathfrak{X}$ be as in the statement of the lemma.
First, we define an auxiliary structure $\mathfrak{X}^{\prime}$ on $X$ with the same signature as $\mathfrak{X}$. For every $k$-ary relation $R^{\mathfrak{B}}$ of $\mathfrak{B}$, the relation $R^{\mathfrak{X}^{\prime}}$ is defined as follows:
[…]
By the definition of full powers, for every $\iota\colon[k]\to[d]$, we have
$\big{(}(x_{\iota(1)},\dots,x_{\iota(k)},y_{k+1}\dots,y_{d}),(x_{1}\dots,x_{k},x_{k+1}\dots,x_{d})\big{)}\in S^{{\mathfrak{X}}^{[d]}}_{\mathrm{id},\iota},$
where $\mathrm{id}$ denotes the identity map on $[k]$. This and the fact that $h$ is a homomorphism yield
$R^{\mathfrak{X}^{\prime}}=\{(x_{1},\dots,x_{k})\in X^{k}\mid h(x_{1},\dots,x_{d})\in R_{\mathrm{id}}^{\ \smash{{\mathfrak{B}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}}}\text{ for all }x_{k+1},\dots,x_{d}\in X\}.$ (2)
Next, we define an auxiliary equivalence relation $\sim$ on $X$ by
$x\sim y:\Leftrightarrow$ $\exists\bar{z}=(z_{1},\dots,z_{d})\in X^{d}\,\exists i_{1},i_{2}\in[d]$
[…]
Here, the containment $\bar{a}\in h(\bar{z})$ is to be understood in the sense that $h(\bar{z})$ represents the orbit of a $d$-tuple. Note that, since $h$ preserves $S_{\iota,\iota^{\prime}}$ for all $\iota,\iota^{\prime}\colon[2]\to[d]$, this is equivalent to
$x\sim y\Leftrightarrow$ $\forall\bar{z}=(z_{1},\dots,z_{d})\in X^{d}\,\forall i_{1},i_{2}\in[d]\colon$
$x=z_{i_{1}}\text{ and }y=z_{i_{2}}\text{ implies }\forall\bar{a}=(a_{1},\dots,a_{d})\in h(\bar{z})\colon a_{i_{1}}=a_{i_{2}}.$ (3)
Claim B.3.
$\sim$ is a relational congruence on $\mathfrak{X}^{\prime}$ w.r.t. all its relations.
Proof.
We must show that, for every $k$-ary relation $R^{\mathfrak{X}^{\prime}}$, every $i\in[k]$, and all $y_{1},\dots,y_{i-1},$ $y_{i+1},\dots,y_{k},x_{1},x_{2}\in X$: if $x_{1}\sim x_{2}$, then
$R^{\mathfrak{X}^{\prime}}(y_{1},\dots,y_{i-1},x_{1},y_{i+1},\dots,y_{k})\Leftrightarrow R^{\mathfrak{X}^{\prime}}(y_{1},\dots,y_{i-1},x_{2},y_{i+1},\dots,y_{k}).$
To this end, define
$\bar{y_{1}}\coloneqq(y_{1},\dots,y_{i-1},x_{1},y_{i+1},\dots,y_{k}),\quad\bar{z_{1}}\coloneqq(y_{1},\dots,y_{i-1},x_{1},y_{i+1},\dots,y_{k},x_{2},\dots,x_{2}),$
$\bar{y_{2}}\coloneqq(y_{1},\dots,y_{i-1},x_{2},y_{i+1},\dots,y_{k}),\quad\bar{z_{2}}\coloneqq(y_{1},\dots,y_{i-1},x_{2},y_{i+1},\dots,y_{k},x_{1},\dots,x_{1}).$
Here, the arity of the tuples $\bar{z}_{1},\bar{z}_{2}$ is $d$. As before, we denote by $\mathrm{id}$ the identity map on $[k]$. Additionally, by $(i\mapsto k+1)$, we denote the map from $[k]$ to $[d]$ sending $i$ to $k+1$ and fixing all other elements. This map is well-defined since all relations are of arity $\leq d-1$. By eq. (3), for all $\bar{a}=(a_{1},\dots,a_{d})\in h(\bar{z}_{1})\cup h(\bar{z}_{2})$, we have $a_{i}=a_{k+1}$. Hence:
$\bar{y_{1}}\in R^{\mathfrak{X}^{\prime}}$ (1)&(2)${\Longleftrightarrow}h(\bar{z_{1}})\in R_{\mathrm{id}}^{\,\smash{{\mathfrak{B}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}}}$
[…]
This is what we had to show. ∎
Finally, we combine the two auxiliary constructions from before by setting $\tilde{\mathfrak{X}}\coloneqq\mathfrak{X}^{\prime}/{\sim}$.
Claim B.4.
$\tilde{\mathfrak{X}}$ embeds into $\mathfrak{B}$.
Proof.
Suppose, on the contrary, that this is not the case. Then there is some bound $\mathfrak{F}$ of $\mathfrak{B}$ that embeds into $\tilde{\mathfrak{X}}$. Without loss of generality, $F=\{[x_{1}]_{\sim},\dots,[x_{n}]_{\sim}\}\subseteq\tilde{X}$. Since all bounds have size $\leq d$, we have $n\leq d$. As before, we shall use the abbreviations $\bar{x}\coloneqq(x_{1},\dots,x_{d})$ and $\bar{a}\coloneqq(a_{1},\dots,a_{d})$. Now, for every $k$-ary relation $R^{\tilde{\mathfrak{X}}}$ and every $\iota\colon[k]\to[n]$, we have
$([x_{\iota(1)}]_{\sim},\dots,[x_{\iota(k)}]_{\sim})\in R^{\tilde{\mathfrak{X}}}$ $\Leftrightarrow\exists x_{1}^{\prime}\in[x_{\iota(1)}]_{\sim},\dots,x_{k}^{\prime}\in[x_{\iota(k)}]_{\sim}\text{ s.t.\ }(x_{1}^{\prime},\dots,x_{k}^{\prime})\in R^{\mathfrak{X}^{\prime}}$
$\Leftrightarrow(x_{\iota(1)},\dots,x_{\iota(k)})\in R^{\mathfrak{X}^{\prime}}$
$\Leftrightarrow\forall x_{n+1},\dots,x_{d}\in X\colon h(\bar{x})\in R_{\iota}^{\smash{{\mathfrak{B}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}}}$
$\Leftrightarrow\forall x_{n+1},\dots,x_{d}\in X\,\forall\bar{a}\in h(\bar{x})\colon\bar{a}\in R_{\iota}^{{\mathfrak{B}}^{[d]}}$
$\Leftrightarrow\forall x_{n+1},\dots,x_{d}\in X\,\forall\bar{a}\in h(\bar{x})\colon(a_{\iota(1)},\dots,a_{\iota(k)})\in R^{\mathfrak{B}}.$
By fixing some $x_{n+1},\dots,x_{d}\in X$ and setting $e([x_{i}]_{\sim})\coloneqq a_{i}$ for $i\leq n$ and some $\bar{a}\in h(\bar{x})$, we have that $e$ is an embedding of $\mathfrak{F}$ into $\mathfrak{B}$, a contradiction. ∎
Let $e^{\prime}$ be an embedding from $\tilde{\mathfrak{X}}$ to $\mathfrak{B}$, which exists by Claim B.4. It follows from Claim B.3 that $e^{\prime}$ induces a homomorphism $g\colon\mathfrak{X}^{\prime}\to\mathfrak{B}$ via $g(x)\coloneqq e^{\prime}([x]_{\sim})$.
Claim B.5.
$h(\bar{x})=g(\bar{x})_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$ for every $\bar{x}\in X^{d}$.
Proof.
Let $i,j\in[d]$ be arbitrary. If every $\bar{y}\in h(\bar{x})$ satisfies $y_{i}=y_{j}$, then $x_{i}\sim x_{j}$ and thus $g(x_{i})=g(x_{j})$. On the other hand, if $g(x_{i})=g(x_{j})$, then $x_{i}\sim x_{j}$ since $e^{\prime}$ is an embedding. Thus, every $\bar{y}\in h(\bar{x})$ satisfies $y_{i}=y_{j}$. Moreover, for all $x_{1},\dots,x_{d}\in X$, every $k$-ary relation $R^{\mathfrak{X}^{\prime}}$, and every $\iota\colon[k]\to[d]$, we have
$h(\bar{x})\in R_{\iota}^{\smash{{\mathfrak{B}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}}}$ $\Leftrightarrow(x_{\iota(1)},\dots,x_{\iota(k)})\in R^{\mathfrak{X}^{\prime}}$
$\Leftrightarrow([x_{\iota(1)}]_{\sim},\dots,[x_{\iota(k)}]_{\sim})\in R^{\tilde{\mathfrak{X}}}$
$\Leftrightarrow(g(x_{\iota(1)}),\dots,g(x_{\iota(k)}))\in R^{\mathfrak{B}}$
$\Leftrightarrow(g(x_{1}),\dots,g(x_{d}))_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}\in R_{\iota}^{\smash{{\mathfrak{B}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}}}.$
Now the statement of the claim follows from the homogeneity of $\mathfrak{B}$. ∎
Claim B.6.
$g$ is a homomorphism from $\mathfrak{X}$ to $\mathfrak{A}$.
Proof.
Let $R\in\tau$ and $(x_{1},\dots,x_{k})\in R^{\mathfrak{X}}$ be arbitrary. Since $h$ is a homomorphism, we have $h(x_{1},\dots,x_{d})\in R_{\mathrm{id}}^{\,\smash{{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}}}$ for all $x_{k+1},\dots,x_{d}\in X$. Since every automorphism of $\mathfrak{B}$ is also an automorphism of $\mathfrak{A}$, it follows from Claim B.5 that $(g(x_{1}),\dots,g(x_{k}))\in R^{\mathfrak{A}}$. ∎
This finishes the proof of the lemma. ∎
Lemma B.7.
Let $\mathfrak{A}$ be a reduct of a finitely bounded homogeneous structure $\mathfrak{B}$ and let $\mathfrak{S}$ be a finite substructure of $\mathfrak{A}$. Suppose that the bounds of $\mathfrak{B}$ have size $\leq d$ and its relations are of arity $\leq d-1$. If $\Sigma$ is a height-1 condition that is satisfied in $\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$, then $\Sigma$ is satisfiable in $\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ modulo $\mathrm{Aut}(\mathfrak{B})$.
Proof.
It is not hard to show that ${(\mathfrak{S}^{n})}^{[d]}\simeq({\mathfrak{S}}^{[d]})^{n}$ (cf. [57, Corollary 22]) for all $d,n\in\mathbb{N}$. Hence, every homomorphism $h\colon({\mathfrak{S}}^{[d]})^{n}\to{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$ is also a homomorphism from ${(\mathfrak{S}^{n})}^{[d]}$ to ${\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$. Suppose that $h$ is a function symbol appearing in $\Sigma$ that is assigned an element of $\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$. By Lemma B.2, there is a homomorphism $g\colon\mathfrak{S}^{n}\to\mathfrak{A}$ such that
$h\begin{pmatrix}\bar{s_{1}}\\ \vdots\\ \bar{s_{d}}\end{pmatrix}=\begin{pmatrix}g(\bar{s_{1}})\\ \vdots\\ g(\bar{s_{d}})\end{pmatrix}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$
for all $\bar{s_{1}},\dots,\bar{s_{d}}\in S^{n}$. For each function symbol in $\Sigma$, we fix such a homomorphism. Suppose now that $h_{1},h_{2}\in\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ witness the satisfaction of the identity
$h_{1}(x_{i_{1}},\dots,x_{i_{n}})\approx h_{2}(x_{i_{n+1}},\dots,x_{i_{n+k}}),$
where $i_{\ell}\leq m$ for some $m\in\mathbb{N}$, and let $g_{1}:\mathfrak{S}^{n}\to\mathfrak{A}$ and $g_{2}:\mathfrak{S}^{k}\to\mathfrak{A}$ be the associated homomorphisms.
Let $\bar{s}_{1},\dots,\bar{s}_{m}$ be arbitrary tuples over $S$ of arity $k$ large enough so that, for every $\bar{s}\in S^{m}$, there exists $\ell\in[k]$ such that $\bar{s}$ equals the $\ell$-th entry in $(\bar{s}_{1},\dots,\bar{s}_{m})$ (viewed as a $k\times m$-matrix). Define the $m$-ary operations $\tilde{g_{1}}$ and $\tilde{g_{2}}$ by $\tilde{g_{1}}(x_{1},\dots,x_{m})\coloneqq g_{1}(x_{i_{1}},\dots,x_{i_{n}})$ and $\tilde{g_{2}}(x_{1},\dots,x_{m})\coloneqq g_{2}(x_{i_{n+1}},\dots,x_{i_{n+k}})$. Then the tuples $\tilde{g_{1}}(\bar{s}_{1},\dots,\bar{s}_{m})$ and $\tilde{g_{2}}(\bar{s}_{1},\dots,\bar{s}_{m})$ satisfy the same equalities and relations w.r.t. $\mathfrak{B}$ on all $d$-tuples. Since all relations of $\mathfrak{B}$ have arity $\leq d$ and $\mathfrak{B}$ is homogeneous, $\tilde{g_{1}}(\bar{s}_{1},\dots,\bar{s}_{m})$ and $\tilde{g_{2}}(\bar{s}_{1},\dots,\bar{s}_{m})$ lie in the same orbit. Thus, there are $\alpha_{1},\alpha_{2}\in\mathrm{Aut}(\mathfrak{B})$ such that
$\alpha_{1}\circ\tilde{g_{1}}(\bar{s}_{1},\dots,\bar{s}_{m})=\alpha_{2}\circ\tilde{g_{2}}(\bar{s}_{1},\dots,\bar{s}_{m}).$
This is equivalent to
$\alpha_{1}\circ g_{1}(s_{i_{1}},\dots,s_{i_{n}})=\alpha_{2}\circ g_{2}(s_{i_{n+1}},\dots,s_{i_{n+k}})$
for all $s_{1},\dots,s_{m}\in S$. Thus, $\Sigma$ is satisfied in $\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ modulo $\mathrm{Aut}(\mathfrak{B})$. ∎
Now we can finally prove Proposition 5.2.
Proof of Proposition 5.2.
Clearly, for every finite substructure $\mathfrak{S}$ of $\mathfrak{A}$, we have ${\mathfrak{S}}^{[d]}\rightarrow{\mathfrak{A}}^{[d]}\rightarrow{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$. Let $S\subseteq A$ with $|S|\geq 3$. Suppose that there is a finite cheese $\mathfrak{D}$ of $({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ that does not pp-construct $\mathfrak{K}_{3}$. Then, by Theorem 3.1, there exists a cyclic operation $f$ in $\mathrm{Pol}(\mathfrak{D})$ of some arity $n$. Take any homomorphisms $g\colon{\mathfrak{S}}^{[d]}\rightarrow\mathfrak{D}$ and $h\colon\mathfrak{D}\rightarrow{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$. Then, by setting $f^{\prime}(x_{1},\dots,x_{n})\coloneqq h\circ f(g(x_{1}),\dots,g(x_{n})),$ we get a cyclic $n$-ary $f^{\prime}\in\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}).$ By Lemma B.7 there is some $n$-ary $g\in\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ that is pseudo-cyclic modulo $\mathrm{Aut}(\mathfrak{B})$. By Lemma B.1, this operation is cyclic. Take pairwise distinct elements $s_{1},s_{2},s_{3}\in S$. We now show that the cyclicity of $g$ contradicts the fact that it preserves $\neq$. To this end, we consider the following three cases:
Case $n\equiv 0\bmod 3$:
$g(s_{1},s_{2},s_{3},\dots,s_{1},s_{2},s_{3})=g(s_{2},s_{3},s_{1},\dots,s_{2},s_{3},s_{1})$
Case $n\equiv 1\bmod 3$:
$g(s_{1},s_{2},s_{3},\dots,s_{1},s_{2},s_{3},s_{1})=g(s_{3},s_{1},s_{2},\dots,s_{3},s_{1},s_{1},s_{2})$
Case $n\equiv 2\bmod 3$:
$g(s_{1},s_{2},s_{3},\dots,s_{1},s_{2},s_{3},s_{1},s_{2})=g(s_{2},s_{3},s_{1},\dots,s_{2},s_{3},s_{1},s_{2},s_{1})$
In all three cases, $g$ does not preserve $\neq$. Hence, the original assumption that such $\mathfrak{D}$ exists was wrong, and consequently $\mathrm{PCSP}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$ is not finitely tractable. ∎
B.3. A full proof of Theorem 2.6
See 2.6
Proof.
[…]
We select a finite substructure $\mathfrak{S}$ of $\mathfrak{A}$ with $|S|\geq 3$ large enough so that every finite structure which has a homomorphism to a substructure of $\mathfrak{A}$ of size $\leq n$ also has a homomorphism to $\mathfrak{S}$ (this can be done since $\mathfrak{A}$ has finite signature). Now let $\widehat{\mathfrak{S}}$ be the substructure of $\widehat{\mathfrak{A}}$ induced by $\{0\}\times S$.
We claim that $\mathfrak{A}^{\prime}\coloneqq{\widehat{\mathfrak{A}}}^{[d]}$, $\mathfrak{B}^{\prime}\coloneqq{\widehat{\mathfrak{B}}}^{[d]}$, $\mathfrak{S}_{1}\coloneqq{\widehat{\mathfrak{S}}}^{[d]}$, and $\mathfrak{S}_{2}\coloneqq{\widehat{\mathfrak{A}}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\widehat{\mathfrak{B}})}}}$ for $d$ as in Proposition 5.2 witness items 17 of the theorem.
Item 1 follows from Theorem 2.2(4) and Proposition 5.1(2).
Item 2 follows from Theorem 2.2 and Proposition 5.1(2).
Item 3 follows from Theorem 2.2(1) and Proposition 5.1(5).
Item 4 follows directly from Theorem 2.2(2), Proposition 5.1(7), and [14, Theorem 1.5].
Items 5 and 6 follow from Proposition 5.2.
For item 7, note that if $\mathfrak{X}$ maps homomorphically to a substructure of $\mathfrak{A}$ of size $\leq n$, then $\mathfrak{X}\rightarrow\mathfrak{S}$. Since the Datalog reduction $\mathcal{I}$ from Theorem 2.2 is just the expansion of a structure by empty relations $\neq$ and $E$, we have $\mathcal{I}(\mathfrak{X})\rightarrow\widehat{\mathfrak{S}}$. Since every homomorphism between two structures clearly induces a homomorphism between their corresponding full powers, we have $\mathcal{J}(\mathfrak{X})={\mathcal{I}(\mathfrak{X})}^{[d]}\rightarrow{\widehat{\mathfrak{S}}}^{[d]}=\mathfrak{S}_{1}$. On the other hand, if $\mathfrak{X}\nrightarrow\mathfrak{A}$, then $\mathcal{I}(\mathfrak{X})\nrightarrow\widehat{\mathfrak{A}}$ by Theorem 2.2. By Lemma B.2, if ${\mathcal{I}(\mathfrak{X})}^{[d]}\rightarrow\mathfrak{S}_{2}$, then also $\mathcal{I}(\mathfrak{X})\rightarrow\widehat{\mathfrak{A}}$, hence $\mathcal{J}(\mathfrak{X})={\mathcal{I}(\mathfrak{X})}^{[d]}\nrightarrow\mathfrak{S}_{2}$. ∎